### v1

# DeltaPhi: Learning Physical Trajectory Residual for PDE Solving

Abstract
Although neural operator networks theoretically approximate any operator mapping, the limited generalization capability prevents them from learning correct physical dynamics when potential data biases exist, particularly in the practical PDE solving scenario where the available data amount is restricted or the resolution is extremely low. To address this issue, we propose and formulate the Physical Trajectory Residual Learning (DeltaPhi), which learns to predict the physical residuals between the pending solved trajectory and a known similar auxiliary trajectory. First, we transform the direct operator mapping between input-output function fields in original training data to residual operator mapping between input function pairs and output function residuals. Next, we learn the surrogate model for the residual operator mapping based on existing neural operator networks. Additionally, we design helpful customized auxiliary inputs for efficient optimization. Through extensive experiments, we conclude that, compared to direct learning, physical residual learning is preferred for PDE solving.
1 Introduction
Due to lacking analytical solutions, solving complex Partial Differential Equations (PDEs), e.g. Navier-Stokes, relies on numerical simulation based on ultra-fine grid division, which consumes expensive computational resources and time. To resolve this, machine learning based PDE solving [1, 2, 3, 4, 5] has sparked extensive research interest in recent years. Notably, the operator learning [6, 2, 1] directly learns the mapping between the input-output function fields, forming a surrogate model for a given parametric PDE. Then the solution for any input functions could be quickly obtained via neural network forward computation.
Despite current operator learning networks [2, 3] theoretically being able to approximate any operator mapping, their generalization capabilities remain limited. First, the networks are prone to overfitting to training resolution while suffering performance degradation on other resolutions [7], especially when the training resolution is extremely low. Second, when available training data is limited, the network may struggle to accurately model the correct physical dynamics due to potential biases in the data. Third, data scarcity and bias are common occurrences in practical PDE solving [8, 9]. Therefore, further investigation is required to enhance the neural operator generalization ability.
In this work, for generalized PDE solving, we introduce the Physical Trajectory Residual Learning (DeltaPhi), which reformulates the direct operator mapping between the input-output function fields as the mapping between input function pairs and output function residuals (residual operator mapping). The residual operator mapping could be learned with any existing neural operators (such as FNO [2], NORM [10], FFNO [11], etc. ), thereby obtaining corresponding residual neural operator. First, we reconstruct the training set from data pairs of direct operator mapping to data pairs of residual operator mapping, via assigning an auxiliary trajectory to every training data. The auxiliary trajectory is a different trajectory sampled from the training set based on similar trajectory retrieval. Next, we optimize the neural operator network with reconstructed training data. By concatenating the pending solved input function and auxiliary trajectory together as input of the network, and adding the auxiliary trajectory solution to the final output of the network via residual connection, we could optimize the network with existing training frameworks. All other setups (including the loss function, trainable weights amount, etc. ) remain consistent with the base operator networks. Compared to previous direct operator learning, the reconstructed residual training labels could be customized with enhanced diversity through selecting appropriate auxiliary samples. Consequently, the restricted diversity and potential distribution bias due to limited training data amount could be mitigated, thereby obtaining more generalized surrogate models.
Residual learning methodology has been extensively studied in the machine learning community [12, 13, 14, 15]. Previous works present various benefits derived from residual learning, such as enhancing generalization capability, handling complex regression mapping, etc. In this work, we introduce this concept into the neural operator learning task, a more complex regression task involving learning the mapping between high-dimension (tens of thousands) fields. The residuals between different but similar physical trajectories are learned, and utilized to solve PDEs. Through the visualization analysis of label distribution, we statistically validate the favorable generalization property of learning the residual operator mapping.
The contributions are summarized as follows:
  • We propose and formulate a novel framework for data-driven PDE solving in restricted data (limited amount, low resolution, biased distribution) scenarios, i.e. the Physical Trajectory Residual Learning (DeltaPhi).
  • We statistically analysis the advantage of residual operator learning and make instantiations based on similar trajectory retrieval and existing neural operator networks.
  • We validate the superiority of residual operator learning through extensive experiments across different PDE problems, different physical domains (both regular and irregular problems), different base models, different training set sizes, and different training resolutions. Additionally, detailed ablation studies are conduct.
2 Related Work
Neural operator learning. Fourier Neural Operator (FNO) [2] utilizes the Fourier Transform based integral operation to efficiently learn the physical dynamics using limited training data. Factorized Fourier Neural Operator (FFNO) [11] factorize the Fourier Transform along each dimension, reducing the number of network weights. Clifford Fourier Neural Operator (CFNO) [16] employs the Clifford Algebra in the network architecture, incorporating geometry prior between multi physical fields. The various neural operators could be simply integrated with the proposed Physical Trajectory Residual Learning.
Other works investigate architecture improvements for different needs, including chaotic systems modeling [17], physical-informed instance-wise finetuning [7], accelerating computation [18, 19, 20], spherical fields processing [21], irregular fields learning [22, 23], non-periodic boundary fields modeling [24], large-scale pretraining [25, 26], etc. In addition, some works explore other network backbones, e.g. improved frequency-spatial domain transformation [27, 28, 29, 30], convolutions [31], graph neural network [32, 6, 33], attention mechanism [34, 23, 35], and diffusion models [36]. Incorporating residual learning with these designs is worth exploring in the future.
Residual learning in previous works. Some works learn residual between different data samples. Similar to ResMem [12], we also utilize the training set as auxiliary memory during inference, enjoying its generalization enhancement property. The difference lies that (1) we memorize the auxiliary labels and learn the residuals, while ResMem memorizes the residuals and learns the labels, (2) our framework is end-to-end trained while ResMem is trained in separate stages. Similar external memory augmented strategies are popular in the Language Modeling community [37, 38].
Other works learn residuals between low resolution and high resolution representation for image super-resolution [39] and deep learning based CFD simulation [39]. Similar residual learning is investigated in time series prediction task  [40, 13]. The primary distinction between these works to ours is that we do not utilize the corresponding low-resolution solutions or previous time-step solutions, but instead, learn the residuals between different trajectories.
3 Methodology
3.1 Preliminary
3.1.1 Direct learning formulation
Direct operator mapping. Previous works [2, 11, 23] formulate the parametric PDE numerical solution as an operator mapping between two infinite dimensional function spaces:
$\mathcal{G}:\mathcal{A}\longrightarrow\mathcal{U},$ (1)
where $\mathcal{A}$ and $\mathcal{U}$ are both banach spaces of functions. For steady-state solving problems, e.g. Darcy Flow, $a(x)\in\mathcal{A}$ represents the diffusion coefficient function, and $u(x)\in\mathcal{U}$ is the corresponding solution function. For time-series forecasting problems, e.g. Navier-Stokes, $a(x)\in\mathcal{A}$ typically denotes the state function of the preceding few steps, while $u(x)\in\mathcal{U}$ represents the function in the subsequent few steps. The learning task is to study a parameterized surrogate model $\mathcal{G}_{\theta}$ for $\mathcal{G}$.
Domain discretization. Following [2, 41], we numerically process the function $a(x)$ and $u(x)$ by taking the discretization of the continuous domain. For regular domain problems, we uniformly discretize the regular field, using standard rectangular grids to capture the physical states. For irregular domain problems, we take the discrete mesh representation of the field.
Optimization problem. Consider the training set $\mathcal{T}$ consisting of $N$ i.i.d. pairs $\{a_{i},u_{i}\}_{i=1}^{N}$, where $a_{i}\in\mathcal{A}$, $u_{i}\in\mathcal{U}$, the objective is to find $\theta$ that minimizes the following cost function:
$\ell=\sum_{i=1}^{N}\mathcal{L}(\mathcal{G}_{\theta}(a_{i}),u_{i}),$ (2)
where $\mathcal{L}$ is a cost function that measures the difference between two functions. During inference, given the input function $a\in\mathcal{A}$, we obtain the corresponding solution $u=\mathcal{G}_{\theta}(a)$.
3.1.2 Neural Operator Network
Prior works [2, 34, 11] have extensively explored neural network architectures for learning the operator mapping $\mathcal{G}_{\theta}$ between function spaces. Among these, the Fourier Neural Operator (FNO)  [2, 1] has been the most commonly applied model across various physical systems. Neural operators are defined as multi-layer networks formulated as follows:
$\mathcal{G}_{\theta}=Q\circ\sigma(W_{l}+\mathcal{K}_{l})\circ\cdots\circ\sigma(W_{i}+\mathcal{K}_{i})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1})\circ P,$ (3)
where $l$ represents the number of layers in the network. The projector $P$ encodes the input function $a_{i}$ into an implicit embedding $v_{0}$. The operators $W_{i}$ and $\mathcal{K}_{i}$ at the $i$-th layer transform the input function $v_{i-1}$ into $v_{i}$. $\sigma$ is an element-wise non-linear operation introduced for learning non-linear mapping between function spaces. $Q$ is the output projector which decodes the last layer’s output $v_{l}(x)$ into the targeted function $u_{i}(x)$.
The components of $\mathcal{G}_{\theta}$ could be differently instantiated across various models. Typically, $P$ and $Q$ are implemented as fully connected layers. In Fourier Neural Operator [2], $\mathcal{K}_{i}$ is instantiated as an integral operator based on Fourier Transform, and $W_{i}$ is a linear fully connected layer. In NORM [10], $\mathcal{K}_{i}$ is implemented based on the Laplacian Transform for processing irregular domains. In addition, $\mathcal{K}_{i}$ can also be a convolution operation, adopting an architecture similar to Resnet [42].
Although direct operator learning networks can theoretically approximate any operator mapping, they encounter limited generalization capabilities in practical applications. First, the networks tend to overfit on trained resolution, and suffer performance drops on high resolution inference. Second, the model is vulnerable to potential bias in training data (especially when the data amount is limited), thus failing to capture the intrinsic physical trajectory dynamics.
3.2 Physical Trajectory Residual Learning
Figure 1: The overall architecture of Physical Trajectory Residual Learning. Given an input trajectory $a_{i}$, we first sample a similar auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$ from the training trajectory set $\mathcal{T}$. Subsequently, the trajectory $a_{i}$ and $(a_{k_{i}},u_{k_{i}})$ are concatenated and fed into the the neural operator $\mathcal{G}_{\theta}^{\Delta}$, producing the predicted trajectory residual. Finally, the predicted solution $\hat{u}_{i}$ is obtained by adding the predicted trajectory residual with the auxiliary solution $u_{k_{i}}$.
3.2.1 Formulation
Different from the aforementioned direct operator learning, we propose to learn the residual operator mapping that captures the differences between pairs of physical trajectories:
$\mathcal{G}^{\Delta}$ $:\mathcal{A}^{2}\longrightarrow\Delta\,\mathcal{U},$ (4)
$\mathcal{A}^{2}$ :$:\{(a_{i},a_{j})\mid a_{i}\in\mathcal{A},a_{j}\in\mathcal{A}\},$
$\Delta\,\mathcal{U}$ :$:\{u_{i}-u_{j}\mid u_{i}\in\mathcal{U},u_{j}\in\mathcal{U}\},$
where $\mathcal{G}^{\Delta}$ represents the residual mapping operator. $\mathcal{A}^{2}$ is the Cartesian product of $\mathcal{A}$ with itself, meaning it consists of all possible pairs of functions from $\mathcal{A}$. $\Delta\,\mathcal{U}$ represents the space of function residuals, which are the differences between pairs of functions from $\mathcal{U}$. $a_{*}$ and $u_{*}$ are functions in the space of $\mathcal{A}$ and $\mathcal{U}$, respectively.
The residual mapping operator $\mathcal{G}^{\Delta}$ maps the input functions pair $(a_{i},a_{j})$ into the corresponding function residual $u_{i}-u_{j}$, i.e. $u_{i}-u_{j}=\mathcal{G}^{\Delta}(a_{i},a_{j})$. This approach leverages the concept of residuals to simplify the learning task. Instead of learning the direct operator mapping from $\mathcal{A}$ to $\mathcal{U}$, we focus on learning the differences between output functions. The transformed training label distribution, i.e. the function residual distribution, could have sufficient diversity by selecting appropriate auxiliary samples, thereby alleviating the training data bias and better handling the unknown function to be solved. We detail the explanation about this in Section 3.2.5.
Both direct operator mapping and residual operator mapping are mappings between Banach spaces of functions. Therefore, we could employ the off-the-shelf neural operator networks [2, 11] as the base network backbone. Neural operator networks are designed to handle mapping between function spaces and are known for their universal approximation capabilities, making them well-suited for residual operator mapping.
3.2.2 Auxiliary Trajectory Sampling
Given an input function (or named as input trajectory) $a_{i}$, we need an auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$ to learn the residual neural operator $\mathcal{G}^{\Delta}$ and infer the corresponding output $u_{i}$ of $a_{i}$. The auxiliary trajectory can be sampled from the existing data, allowing us to leverage prior knowledge to improve the generalization capabilities of the model. By selecting an appropriate auxiliary trajectory, we can simplify the computation and enhance the robustness of the residual operator mapping.
Given the training set $\mathcal{T}=\{a_{i},u_{i}\}_{i=1}^{N}$, each data pair $(a_{i},u_{i})$ is associated with an auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$, sampled from $\mathcal{T}$. The index $k_{i}$ of the auxiliary trajectory is defined by:
$k_{i}=\mathcal{F}^{ret}(a_{i},\mathcal{T}),\,k_{i}\in\{1,2,...,N\},\,k_{i}\neq i,$ (5)
where $\mathcal{F}^{ret}$ is the sampling function. A straightforward implementation of $\mathcal{F}^{ret}$ could be random sampling from $\{1,2,...,i-1,i+1,...N\}$. Using the calculated $k_{i}$, we transform the origin training set $\mathcal{T}$ to a residual training set $\mathcal{T}^{res}=\{(a_{i},a_{k_{i}}),u_{i}-u_{k_{i}}\}_{i=1}^{N}$. Below we present our implementation of $\mathcal{F}^{ret}$ in detail.
Training stage. During training, give an input function $a_{i}$, we randomly sample a trajectory $a_{j}$ from the set of trajectories with top-k similarity scores to $a_{i}$. The sampling function $\mathcal{F}^{ret}$ is defined as:
$\mathcal{F}^{ret}(a_{i},\mathcal{T})$ $=\texttt{RandSample}(\texttt{argtopK}_{j,j\neq i}\,\texttt{Sim}(a_{i},a_{j})),$ (6)
$\texttt{Sim}(a_{i},a_{j})$ $=1-{a_{i}\cdot a_{j}\over\|a_{i}\|\|a_{j}\|},$
where $\texttt{RandSample}(\cdot)$ represents randomly sampling one element from the given set. $K$ is the sampling range, a hyperparameter that could be adjusted for different PDEs. For a fair comparison, except for specific statements, we set $K=20$ for all experiments in this work.
Inference stage. During inference, given the input function $a^{test}$, we select the most similar auxiliary trajectory from the training set:
$\mathcal{F}^{ret}(a^{test},\mathcal{T})=\texttt{argmax}_{j}\,\texttt{Sim}(a^{test},a_{j}).$ (7)
We empirically found the model performance is robust to the choice of auxiliary trajectories. In Section A.3, we experiment with randomly sampled auxiliary trajectories from the top ten similarity scores and find the standard deviation across five repeating evaluations is quite small. This indicates that the residual neural operator effectively learns the residual operator mapping for the same $a_{i}$ with varied auxiliary trajectories. We further present the trajectory sampling for varied resolution fields in Section C.3. Section C.4 provides the cost analysis of the trajectory sampling.
3.2.3 Auxiliary Trajectory Integration
As aforementioned, the residual operator $\mathcal{G}^{\Delta}$ requires additional auxiliary function $a_{k_{i}}$ as input. In specific implementation, $a_{i}$ and $a_{k_{i}}$ could be straightly concatenated along their channel dimension:
$a_{input}=a_{i}\oplus a_{k_{i}},$ (8)
where $\oplus$ represents the concatenation operation of two vectors. Then the concatenated vector $a_{input}$ is feeded into the neural operator network:
$v_{0}=P(a_{input}),$ (9)
where $P$ is the fully connected encoder in the operator network (Equation 3), and $v_{0}$ represents the first latent function in the hidden space.
Integration of auxiliary trajectory with variant resolution. In cases where the auxiliary input function $a_{k_{i}}$ has a different resolution from the input function $a_{i}$, we address this discrepancy via field interpolation techniques. We present the details in Section C.2.
Integration of auxiliary trajectory with additional related information. The residual operator mapping can be challenging to learn due to the variability of output residuals for different auxiliary samples. To mitigate such difficulty, we introduce additional information related to $a_{k_{i}}$ as inputs for residual neural operators:
  • For time-series PDE solving, e.g. Navier-Stokes, we only input the last few time steps of $a_{k_{i}}$.
  • We use the corresponding $u_{k_{i}}(x)$ of $a_{k_{i}}(x)$ as the input.
  • We utilize calculated relation metrics, e.g. similarity between $a_{i}$ and $a_{k_{i}}$, as inputs.
Our experimental results (Section A.2) indicate that this customized input strategy indeed improves the performance of residual neural operators.
3.2.4 Residual Operator Learning
To train the parameterized surrogate model $\mathcal{G}_{\theta}^{\Delta}$ of $\mathcal{G}^{\Delta}$ using the residual training set $\mathcal{T}^{res}$, we formulate the following optimization problem to determine the optimal parameter by minimizing the cost function:
$\ell=\sum_{i=1}^{N}\mathcal{L}(\mathcal{G}_{\theta}^{\Delta}(a_{i},a_{k_{i}}),u_{i}-u_{k_{i}}),$ (10)
where $\mathcal{L}$ is the cost function for assessing the discrepancy between function residuals.
The optimization of the physical trajectory residual depicted in Equation 10 can be effectively implemented using existing learning frameworks [2, 11]. This involves introducing a trajectory residual connection on the explicit output of the base operator network. Specifically, we add the auxiliary solution $u_{k_{i}}$ to the predicted residual from the final output of the operator network:
$\hat{u}_{i}=Q(v_{l})+u_{k_{i}},$ (11)
where $Q$ is the last fully connected projector of the operator network, $v_{l}$ is the last latent function in hidden space produced by $\sigma(W_{l}+\mathcal{K}_{l})$ shown in Equation 3.
During inference, given input function $a^{test}$, we select the auxiliary trajectory $(a_{k_{t}}(x),u_{k_{t}}(x))$ from the training set $\mathcal{T}$ using the sampling function $\mathcal{F}^{ret}$.
3.2.5 Discussion
Compared to direct operator learning, residual operator learning presents the following advantages.
Mitigation of distribution bias. Residual operator learning allows us to control the training label distribution, thereby mitigating the data distribution bias. We posit that the similarity between input functions and the corresponding output functions’ similarity are positively correlated. This correlation is validated empirically in non-chaotic systems, as demonstrated in Darcy Flow experiments (Figure 4). Consequently, the distribution of label residuals can be effectively controlled by adjusting the selection range ($K$ in Equation 6) of the sampling function $\mathcal{F}^{ret}$. This adjustment allows the construction of sufficiently diverse training labels, thereby alleviating distribution bias in restricted data scenarios and enhancing the generalization capabilities of neural operators.
Prevent overfitting. In scenarios with limited training data, neural networks tend to memorize the training set, leading to significant overfitting. Residual learning effectively addresses this issue by modifying the training labels through the use of various auxiliary samples in each training epoch, thus reducing the risk of overfitting.
Incorporation of explicit physical priors. The auxiliary sample acts as a non-parametric physical prior. As shown in previous work [12], involving external memory components benefits the model’s generalization capability. Moreover, the explicit external memory allows specific needs during inference, such as handling boundary conditions and high resolution samples.
4 Experiment
This section provides the experiments on various PDEs, resolution generalization problem, and statistical analysis of residual operator learning.  Appendix A presents additional ablation studies.
4.1 Residual Learning vs. Direct Learning on Various PDEs
This section validates residual operator learning on various PDEs. The experiment covers both irregular domain problems and regular domain problems. In addition, we test the performance improvement under variant training data amount. To ensure fair comparisons, all baselines and our counterpart models are compared with exactly same hyperparameters, trainable weight amount, optimizer configuration, and training step.
Irregular domain problems. We validate the effectiveness of the proposed approach on irregular domain problems. The experimental setup is exactly same as [10]. We experiment with five problems covering both 2D and 3D scenarios, i.e. Irregular Darcy Flow (2D), Pipe Turbulence (2D), Heat Transfer (3D), Composite (3D), and Blood Flow (3D), all of which are defined on irregular physical domains and processed as a set of discrete triangle meshes. The baselines contain GraphSAGE [43], DeepOnet [3], POD-DeepOnet [44], FNO [2] and NORM [10]. We implement the residual learning version of NORM and compare it with the direct learning model under exactly same setup. Appendix B presents more experimental details about the datasets and baselines.
We report the results in Table 1. The residual neural operator NORM (DeltaPhi) consistently outperforms all baselines on all irregular domain problems. The enhancement is particularly significant on Pipe Turbulence and Heat Transfer problem, with 40% - 50% relative improvement (the row "Relative Gain" in Table 1) over original NORM, i.e. the NORM (Previous). These results indicate that the proposed residual operator learning works well on irregular domain problems.
Figure 2: Prediction error visualization on irregular domain problems.
In addition, we visualize the prediction error in Figure 2. The error value is calculated as the absolute difference between the predicted function $\hat{u}$ and ground-truth function $u^{GT}$, i.e. $|\hat{u}-u^{GT}|$. Compared to previous direct learning methods, DeltaPhi significantly reduces the error scale.
Table 1: Relative error comparison of operator learning on irregular domains.
Model
Irregular Darcy
(Train Size=1000)
Pipe Turbulence
(Train Size=300)
Heat Transfer
(Train Size=100)
Composite
(Train Size=400)
Blood Flow
(Train Size=400)
GraphSAGE 6.73e-2 2.36e-1 - 2.09e-1 -
DeepOnet 1.36e-2 9.36e-2 7.20e-4 1.88e-2 8.93e-1
POD-DeepOnet 1.30e-2 2.59e-2 5.70e-4 1.44e-2 3.74e-1
FNO 3.83e-2 3.80e-2 - - -
NORM (Previous) 1.05e-2 1.01e-2 2.70e-4 9.99e-3 4.82e-2
NORM (DeltaPhi) 9.75e-3 5.96e-3 1.35e-4 9.18e-3 4.29e-2
Relative Gain 7.14% $\uparrow$ 40.99% $\uparrow$ 50.00% $\uparrow$ 8.11% $\uparrow$ 11.00% $\uparrow$
Regular domain problems. We conduct the experiment on two regular domain problems i.e. Darcy Flow and Navier-Stokes, following [2]. The training data amount is 100. We validate DeltaPhi on different neural operators, including FNO [2], FFNO [11], CFNO [16], GNOT [23], Galerkin [34], MiOnet [45] and Resnet [42]. The performance is measured with relative error. Appendix B presents more experimental details.
Table 2 presents the experimental results. On both Darcy Flow and Navier-Stokes Equations, the proposed residual operator learning (DeltaPhi) consistently improves performance across different base models. The results validate the effectiveness of DeltaPhi on regular domain problems. However, we observe that the performance improvement of the model on the Navier-Stokes Equation is relatively marginal. We conjecture this is due to the complexity of the fluid dynamic with the viscosity $\upsilon=1e-5$, whereby the correlation between input and output functions with long time intervals is diminished, i.e. similar initial functions $a(x)$ do not necessarily correspond to sufficiently similar output functions $u(x)$. This is a limitation of the current implementation that does not affect the conclusion of this work. It could be alleviated with additional designs such as performing auxiliary trajectory retrieval at each step of time-series prediction to mitigate the decay in correlation between input and output functions.
Table 2: Relative error comparison on regular Darcy Flow and Navier-Stokes Equation.
Model Darcy Flow Navier-Stokes Equation
Direct Learning
Residual Learning
Relative Gain
Direct Learning
Residual Learning
Relative Gain
FNO 3.70e-2 3.31e-2 10.54% $\uparrow$ 2.24e-1 2.13e-1 4.86% $\uparrow$
FFNO 5.22e-2 2.93e-2 43.76% $\uparrow$ 2.40e-1 2.20e-1 8.54% $\uparrow$
CFNO 4.79e-2 3.15e-2 34.23% $\uparrow$ 3.51e-1 3.35e-1 4.56% $\uparrow$
GNOT 6.74e-2 5.50e-2 18.39% $\uparrow$ 4.16e-1 4.06e-1 2.33% $\uparrow$
Galerkin 6.78e-2 6.54e-2 3.60% $\uparrow$ 3.60e-1 3.54e-1 1.86% $\uparrow$
MiOnet 8.61e-2 8.22e-2 4.53% $\uparrow$ 4.79e-1 4.61e-1 3.80% $\uparrow$
Resnet 1.25e-1 1.07e-1 14.14% $\uparrow$ 4.39e-1 4.30e-1 1.87% $\uparrow$
Table 3: Relative error comparison on different training scales.
Model Different Train Size
100
300
500
700
900
FNO (Previous) 3.70e-2 1.40e-2 1.02e-2 8.37e-3 7.39e-3
FNO (DeltaPhi) 3.31e-2 1.34e-2 9.64e-3 8.06e-3 7.18e-3
Relative Gain 10.54% $\uparrow$ 4.08% $\uparrow$ 5.49% $\uparrow$ 3.74% $\uparrow$ 2.83% $\uparrow$
FFNO (Previous) 5.22e-2 1.69e-2 9.73e-3 7.48e-3 6.16e-3
FFNO (DeltaPhi) 2.93e-2 1.11e-2 7.20e-3 6.22e-3 5.30e-3
Relative Gain 43.76% $\uparrow$ 34.47% $\uparrow$ 25.98% $\uparrow$ 16.79% $\uparrow$ 14.02% $\uparrow$
Resnet (Previous) 1.25e-1 1.22e-1 1.21e-1 1.21e-1 1.21e-1
Resnet (DeltaPhi) 1.07e-1 9.42e-2 9.08e-2 8.63e-2 8.31e-2
Relative Gain 14.14% $\uparrow$ 22.74% $\uparrow$ 25.14% $\uparrow$ 28.92% $\uparrow$ 31.51% $\uparrow$
Different training scales. We test the residual operator learning on different training set sizes. The experiment is conducted on Darcy Flow problems, with three base models FNO, FFNO, and Resnet. Specifically, we train all models with five training data scales 100, 300, 500, 700, and 900 respectively, and evaluate their performance on the additional same 100 testing samples.
The experimental results are shown in Table 3. On different training set sizes, DeltaPhi improves the prediction performances of neural operators. As the number of available training data decreases, the relative improvement percent (the row "Relative Gain") consistently increases on FNO and FFNO. In addition, several residual neural operators even outperform direct neural operators using less training data amount, e.g. FFNO (DeltaPhi) with training size 500 outperforms FFNO (Previous) with training size 700, Resnet (DeltaPhi) with training size 700 outperforms Resnet (Previous) with training size 900. The experimental results empirically demonstrate the effectiveness of DeltaPhi across different training set scales.
4.2 Residual Learning vs. Direct Learning on Resolution Generalization Problem
This section presents the experimental results of training-free resolution generalization.
Resolution generalization problem. Fourier neural operator [2] enables zero-shot resolution generalization inference. However, when the training data grid is excessively coarse, the inference performance over high resolution data drops a lot. The performance degeneration is more serious when learning operator mapping between fields with weaker spatial continuity such as Darcy Flow.
Setup. We conduct the training-free resolution generalization experiment on Darcy Flow [2]. Specifically, we train the models with low resolutions $85\times 85$, $43\times 43$, $31\times 31$, and $22\times 2$ 22 respectively, then evaluate them using high resolution $421\times 421$ test data. Both FNO [2] and FFNO [11] are compared as the base models. The training set scale includes 100 and 900 trajectories.
Table 4: Relative error comparison on zero-shot resolution generalization.
Model Train Size=100 Train Size=900
$\mathbb{85\times 85}$
$\mathbb{43\times 43}$
$\mathbb{31\times 31}$
$\mathbb{22\times 22}$
$\mathbb{85\times 85}$
$\mathbb{43\times 43}$
$\mathbb{31\times 31}$
$\mathbb{22\times 22}$
FNO (Previous) 7.29e-2 1.19e-1 1.49e-1 1.70e-1 6.73e-2 1.11e-1 1.34e-1 1.76e-1
FNO (DeltaPhi) 6.91e-2 1.13e-1 1.33e-1 1.53e-1 4.91e-2 9.37e-2 1.15e-1 1.57e-1
Relative Gain 5.13% $\uparrow$ 5.06% $\uparrow$ 10.43% $\uparrow$ 10.08% $\uparrow$ 27.04% $\uparrow$ 15.65% $\uparrow$ 14.29% $\uparrow$ 10.31% $\uparrow$
FFNO (Previous) 6.64e-2 1.04e-1 1.24e-1 1.43e-1 4.89e-2 1.01e-1 1.27e-1 1.49e-1
FFNO (DeltaPhi) 6.34e-2 9.90e-2 1.18e-1 1.36e-1 4.42e-2 9.00e-2 1.12e-1 1.35e-1
Relative Gain 4.53% $\uparrow$ 5.06% $\uparrow$ 4.84% $\uparrow$ 5.09% $\uparrow$ 9.48% $\uparrow$ 10.72% $\uparrow$ 11.94% $\uparrow$ 9.29% $\uparrow$
Figure 3: Training curve comparison. "DP-*" denotes the proposed residual learning version of base models.
Performance comparison. Table 4 presents the resolution generalization results. Compared to FNO [2], FFNO [11] shows prefer generalization ability for its less number of network weights. In all experimental settings, the proposed physical residual models (tagged with "DeltaPhi") evidently outperform their counterpart base models. The performance improvement of residual operator learning is more noticeable when the training set size is 900. In this setting, FNO (DeltaPhi) managed to match, or even surpass FFNO [11] on resolution $85\times 85$, $43\times 43$, and $22\times 2$ 22, despite using the relatively weaker base network. The results empirically draw the conclusion that physical residual learning improves zero-shot resolution generalization performance.
Training curve comparison. We also report the test loss curve during training in  Figure 3. The training data is 900 and the training resolution is $85\times 85$. As the training proceeds, the relative error on resolution $85\times 85$ data continues to decline. However, after some epochs, the loss on resolution $421\times 421$ gradually increases due to overfitting. The proposed physical residual operator (named with DP-FNO) significantly mitigates this overfitting tendency of the base model (named with FNO).
4.3 Statistical Analysis of Residual Operator Learning
This section further validates the generalization property of DeltaPhi through statistical analysis.
Similarity between output functions. We statistically validate the assumption (mentioned in Section 3.2.5) that the similarities between the input function pair and the corresponding output function pair are positively related. This means that for input function $a^{test}(x)$, the higher similarity score of the auxiliary sample $(a_{k_{t}}(x),u_{k_{t}}(x))$, the lower difference between the targeted solution function $u^{test}(x)$ and auxiliary solution function $u_{k_{t}}(x)$.
Figure 4: The variation curve of the normalized distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as retrieval similarity rank increases. Different color corresponds to different retrieval set sizes.
In Figure 4, we visualize the variation curve of the normalized distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as the similarity rank of $a_{k_{t}}(x)$ increases on Darcy Flow. (Section D.1 presents the visualization details.) The normalized distance consistently increases as the similarity rank increases across different retrieval set size. This result concludes that the similarity between output function pair is positively correlated with the similarity between input function pair, thus the training label distribution could be effectively controlled for alleviating the training label distribution bias.
Training label distribution comparison. Through training label distribution visualization, we statistically validate the enhanced generalization property of DeltaPhi on Darcy Flow. We process both the training dataset and the testing dataset with the following steps. First, the high-dimensional labels (output function) are compressed into 2 dimensions. Then we plot the 2D point of every label and the circles of standard deviation and range for all samples. Section D.2 presents more details.
Figure 5: Label distribution visualization on Darcy Flow. The points represent dimension-reduced labels ($u_{*}(x)$ for direct learning, $u_{*}(x)-u_{k_{*}}(x)$ for residual learning) through Principle Component Analysis. The ellipses with dotted lines and solid lines represent the standard deviation and range, respectively. Green and red color correspond to training set and testing set, respectively.
Figure 5 presents the visualization results. (1) As depicted in the left part of Figure 5, in direct operator learning the restricted amount of training data leads to the limited diversity of training labels. Consequently, it is challenging to comprehensively cover the distribution of testing labels with training labels. We refer to this situation as training label distribution bias. In such scenarios, the trained neural operator models struggles with managing certain solutions to be solved. (2) As depicted in the right part of Figure 5, compared to direct operator learning, the testing label distribution becomes more concentrated, and the training labels distribution becomes more dispersed in residual operator learning. The distribution bias between training labels and testing labels is mitigated as mentioned in Section 3.2.5.
5 Conclusion
This work proposes to learn physical trajectory residuals for PDE solving based on existing neural operator networks. Extensive experiments substantiate that learning residual operator mapping is preferred compared to previous learning direct operator mapping between input-output function fields. We hope the introduced method provides insights for future development of machine learning based PDE solving or other high-dimensional regression problems. In future work, we will explore employing DeltaPhi in practical scenarios of data deficiency and bias, e.g. realistic physical phenomenon forecasting, or high-fidelity physics dynamic modeling.
6 Limitation
Despite the comprehensive validation of the superiority of Physical Residual Operator Learning through an extensive range of experiments, we recognize some inherent limitations that, nevertheless, do not impact the solidity of our conclusions. First, when the correlation between the input and output functions of PDEs is weak, such as in chaotic systems, the performance enhancement of the current implementation may be constrained. Second, compared to direct operator learning, the proposed residual operator learning increases the learning difficulty, thus requiring the base model to have sufficient representation capability. Although beyond the scope of this work, these aspects warrant further exploration in future studies.
Appendix A Ablation Studies
A.1 Influence of Similar Trajectory Retrieval
In this section, we test the impact of different retrieval ranges $K$ presented in Equation 6 under various training data scales.
Setup. The experimental PDEs are Darcy Flow [2] and Navier-Stokes [2]. We randomly sample 100, 300, 500, 700, and 900 data for training, and evaluate the trained models on an additional 200 test data with the same resolution. FNO [2] and Resnet [42] are included as the base model for comparison. Besides, we experiment with varying retrieval range $K$ on different training data scales and evaluate the performance without utilizing Similar Trajectory Retrieval.
(a) Darcy FLow
(b) Navier-Stokes
Figure 6: Influence of similar trajectory retrieval range $K$ (presented in Equation 6). "DP-*" denotes the proposed residual learning version of each base model.
Performance Comparison. The experimental results are shown in  Figure 6. In most settings, the physical trajectory residual learning framework DP-FNO and DP-Resnet outperform their corresponding baselines FNO [2] and Resnet [42] respectively. In the small data regime, this performance enhancement is particularly evident. Additionally, on Darcy Flow in Figure 6(a), as the amount of training data increases, the performance increasing of the weak base model Resnet [42] gradually converges, while the corresponding residual operator DP-Resnet continually improves. Same as Table 3, the experimental results validate that the proposed residual operator learning performs better across various training set scales, especially on small data scenarios.
Influence of Similar Trajectory Retrieval.  Figure 6 presents the physical residual learning with varying similar trajectory retrieval range $K$, as well as the simply random sampling without considering similarity (could also be seen as setting $K$ as the training set size). On various settings, as the sampling similarity range $K$ increases, the overall trend is that the model’s performance first rises and then falls. This suggests that setting an appropriate range $K$ is of significant importance. Another point worth noting is that, in nearly all settings, the model’s performance appears to be the poorest when similarity-based retrieval is removed. This demonstrates the crucial role of Similar Trajectory Retrieval. These experimental results lead to the conclusion that Similar Trajectory Retrieval plays an important role in the physical residual learning framework.
A.2 Customized Auxiliary Input Ablation
This section conducts ablation experiments on the customized auxiliary input.
Table 5: Ablation of customized auxiliary input on Navier-Stokes. Bold font, underline and wavyline respectively denote the best, second best, and inferior results.
Customized Auxiliary Input Relative Error
$a_{k_{i}}(x)$ $u_{k_{i}}(x)$ $Score_{i,{k_{i}}}$
All 2.14e-1
2.20e-1
Partial 2.13e-1
Partial 2.13e-1
Partial 2.34e-1
Partial 2.37e-1
Setup. We evaluate the influence of customized auxiliary input on Navier-Stokes. The training trajectory number is 100. FNO [2] is used as the base model. We test 3 types input of $a_{k_{i}}(x)$, including complete $a_{k_{i}}(x)$, empty $a_{k_{i}}(x)$ and partial (last 3 trajectory steps) $a_{k_{i}}(x)$. In addition, we experiment with the inclusion or omission of $u_{k_{i}}(x)$, as well as the addition or exclusion of $Score_{i,k_{i}}$.
Customized Auxiliary Input Influence. The results are shown in Table 5. The first three rows indicate that utilizing partial $a_{k_{i}}(x)$ yields the best results for $a_{k_{i}}(x)$. When comparing the third row with the fourth row, or the fifth row with the sixth row, it’s evident that incorporating $Score_{i,k_{i}}$ is also meaningful. Additionally, when comparing the third row with the fifth row, as well as the fourth row with the sixth row, the inclusion of $u_{k_{i}}(x)$ proves to be extremely crucial. The above findings suggest that the utilization of customized input is beneficial, thereby confirming the conclusion that proper customized input could improve performance.
A.3 Auxiliary Trajectory Selection during Inference
Table 6: Results of repeat Evaluation using random auxiliary sample with Top-10 similarity score.
Relative Error of FNO (DeltaPhi)
Infer 1 3.3441876e-2
Infer 2 3.3364178e-2
Infer 3 3.3501464e-2
Infer 4 3.3489122e-2
Infer 5 3.3489122e-2
STD 5.40649e-5
As described in Section 3.2.2, we greedily select take the auxiliary sample with the top similarity score during inference. To validate the robustness of the trained residual neural operator on different auxiliary samples, we experiment with randomly sampled auxiliary trajectories from the top-10 similarity scores. The experiment is conducted on Darcy Flow with FNO as the base model.
Table 6 presents the relative error of every experimental run and the standard deviation (STD) across five repeated experiments. The low standard deviation value 5.41e-5 indicates that the trained residual neural operators are not sensitive to the selection of auxiliary samples. This concludes that the neural operator effectively learns the proposed residual operator mapping.
A.4 Different Similarity Function
Table 7: Results of DeltaPhi-FNO using different similarity function.
Similarity Functions Relative Error
Cosine Similarity 3.31e-2
Euclidean Distance 3.31e-2
Manhattan Distance 3.28e-2
We use Cosine Similarity for its two advantages:
  • Unaffected by the magnitude of the numerical value. Thus it could be conveniently applied to various PDEs without considering the numerical scale distribution of physical fields.
  • Taking normalized values in $[0,1]$. We could simply use it as an additional network input. As Table 5 shows, incorporating the similarity scores improves the model’s performance on the Navier-Stokes equation.
To explore more retrieval metrics for scientific machine learning, we experiment with other discrepancy metrics (all the distance values are calculated after normalization), as shown in Table 7. The results indicate that different distance functions yield subtle impacts. Euclidean Distance and Manhattan Distance are also appropriate choices. Additionally, inspired by the great success of the NLP community, conducting retrieval in the latent space may yield superior results. We will explore this in the next step.
Appendix B Experimental Detail
B.1 Dataset
B.1.1 Regular Domain
Darcy Flow. Darcy Flow is a steady-state solving problem. We conduct the experiment on Darcy Flow using the same dataset as [2], consisting of $421\times 421$ resolution fields with Dirichlet boundary. The low resolution data e.g. $22\times 2$ are obtained via uniformly downsampling operation.
Navier-Stokes. Navier-Stokes is a challenging time-series solving problem. We use the public dataset from  [2]. The viscosity of trajectories is $51e-$ 5 ($Re=2000$). The spatial resolution is $64\times 64$. The input and output time step length are both $10$.
B.1.2 Irregular Domain
For these irregular domain problems, we take the exactly same setting with NORM, more details about the dataset can be found in [10].
Irregular Darcy. The irregular Darcy problem is to solve the Darcy Flow equation defined on an irregular domain. The input function $a(x)$ represents the diffusion coefficient field and the output function $u(x)$ is the corresponding pressure field. The domain is represented with a set of triangle meshes consisting of 2290 nodes. Following  [10], we train the neural operators with 1000 data and test them on additional 200 trajectories.
Pipe Turbulence. Pipe Turbulence is a dynamic fluid system described by the Navier-Stokes equation. The computational domain is an irregular pipe shape represented as 2673 triangle mesh nodes. In this problem, the neural operator is required to predict the velocity field in the next frame given the previous velocity field. Same as  [10], 300 trajectories are employed for training, and 100 data are used for evaluation in our experiments.
Heat Transfer. In the Heat Transfer problem, the energy transfer phenomena due to temperature difference is studied. The system evolves under the governing law described by the Heat equation. The neural operator is optimized to predict the 3-dimensional temperature field in 3 seconds given the initial boundary temperature state. The output physical domain is represented by triangle meshes with 7199 nodes. We use 100 data for training and the rest of 100 data for evaluation.
Composite. Composite problem is to predict the deformation field in high-temperature stimulation, which is greatly significant for composites manufacturing. The learned operator is expected to predict the deformation field given the input temperature field. Following [10] The studied geometry in this work is an air-intake structural part of a jet, which is composed of 8232 nodes. The training data size is 400 and the test data size is 100.
Blood Flow. This problem aims to predict blood flow in the aorta, which contains 1 inlet and 5 outlets. The blood flow is simulated as a homogeneous Newtonian fluid. The computational domain is completely irregular and represented by a set of triangle meshes comprising 1656 nodes. The simulated temporal length is 1.21 seconds with the temporal step length of 0.01 seconds. The neural operator predicts the velocity field at different times given the velocity boundary at the inlet and pressure boundary at the outlet. In this problem, we have 400 data for training and 100 data for testing, same as  [10].
B.2 Base Model
Fourier Neural Operator (FNO) [2]. Fourier Neural Operator (FNO) [2] utilizes the Fourier Transform based integral operation to implement the neural operator kernel $\mathcal{K}_{i}$. We implement the FNO with the officially published code under the up-to-date deep learning framework. All model hyperparameters except the hidden channel width ($64$ in our experiment, $32$ in the original implementation) are kept the same as in the original manuscript. The incremental hidden channels enhance the model’s representation capability, obtaining consistent performance improvement in all settings. The optimization setup (Adam [46] optimizer with initial learning rate $0.001$ and weight decay $0.0001$, StepLR scheduler with step size 100 and $\gamma=0.5$) is consistent with official implementation.
Factorized Fourier Neural Operator (FFNO) [11]. Factorized Fourier Neural Operator (FFNO) [11] conducts the Fourier Transform along every dimension independently, reducing the model size and enabling deep layer optimization. We implement FFNO via the independent Fourier Transform along each dimension, the improved residual connections, and the FeedForward-based encoder-decoder. All model hyperparameters and optimization setups are consistent with FNO.
Clifford Fourier Neural Operator (CFNO) [16]. Clifford Fourier Neural Operator (CFNO) [16] employs the Clifford Algebra in the neural network architecture, incorporating geometry prior between multiple physical fields. We employ the official implementation for the model architecture. The signature is set as $(-1,-1)$ and we pad lacked physical channels with zero. All hyperparameters except the channel width (32 in CFNO) remain consistent with FNO. The training setup (including the optimizer and scheduler) is the same as FNO.
Galerkin [34]. Galerkin proposes to learn the neural operator with linear attention. We employ the same model architecture as the original work. The network depth is 4, the head number in the attention layer is set as 2, and the latent dimension in the feedforward layer is 256. We use the Galerkin attention mechanism and set the dropout value as 0.05. In our experiment, we use the same training setup as FNO.
GNOT [23]. GNOT is also an attention-based neural operator structure, which introduces the heterogeneous normalized attention layer to handling various inputs such as multiple system inputs and irregular domain meshes. We employ the official architecture implementation. The network depth is set as 4, the hidden dimension is 64, and the activation function is GeLU. The training setup is the same as FNO.
MiOnet [45]. MiOnet [45] is an enhanced version of DeepOnet [3]. It processes multiple inputs by utilizing multiple branch networks and then merging all branch outputs with several fully connected layers. In our implementation, the depth of all branch layers, trunk layer, and merging layer are set as 4. The latent dimension is set as 128. We take the same training setup with FNO.
Resnet [42]. Resnet [42] is a classical network architecture in image processing. Although it fails in zero-shot resolution generalization inference for learning infinite dimension operator mapping, the powerful capability of capturing local details deserves attention. We implement Resnet by simply replacing the spectral convolution in FNO with $3\times 3$ spatial convolution operation. All model hyperparameters and optimization setups remain consistent with FNO.
B.3 Implementation Details
Except for specific statements, the experimental details are as follows:
  • Retrieval: For training, the auxiliary trajectory retrieval range $K$ is set as 20. During inference, the auxiliary trajectory with the most similarity with $a_{i}$ is retrieved.
  • Input: The customized input function is partial (last 3 channels) $a_{k_{i}}$ and complete $u_{k_{i}}$.
  • Optimizing: We use the Adam [46] optimizer with initial learning rate $0.001$ and weight decay $0.0001$. The StepLR scheduler with step size 100 and $\gamma=0.5$ is used to control the learning rate. The batch size is set as 8. All models are trained for 500 epochs. All experiments could be conducted on a single NVIDIA GeForce RTX 4090 device.
B.4 Metric
The evaluated metric is Relative L2 Error, defined as:
$L2={1\over N}\sum_{i=1}^{N}{\|\hat{u}_{i}-u_{i}\|_{2}\over\|u_{i}\|_{2}},$ (12)
where $\hat{u}_{i}$ and $u_{i}$ are predicted trajectory and real trajectory respectively.
Appendix C Additional Method Detail
C.1 Selection of $K$ Value
The optimal value of $K$ is hard to determine. It depends on the generalization problem type (limited training data problem and resolution generalization problem in this work), the base models, etc. Empirically, the larger the value of K, the better the model’s generalization range, but the learning difficulty also increases. A proper value of $K$ should ensure enough diversity of training residuals for generalizing to the targeted pending solved physical trajectories. Thus, it is advisable to set a larger K value when the discrepancy between the training set and test set is evident, and the model’s representation ability is considerably strong. Through extensive experiments, we find $K=20$ is an applicable value for the focused generalization problem of this work, ie. the limited amount and resolution of training data.
For more challenging generalization problems (eg. serious data bias problems in realistic scenarios), we provide an empirical selection algorithm for the "initial value" decision of $K$ (the "final value" of $K$ is also related to the base model’s capability). Given the training set (noted as $\{a^{train}_{i}\}_{i=1}^{N_{train}}$), suppose the input functions (noted as ${\{a^{test}_{i}\}}_{i=1}^{N_{test}}$) of pending solved physical trajectories is available. The initial value of $K$ could be calculated with the following steps: (1) Retrieve the most similar trajectory (noted as $\{a^{test}_{k_{i}}\}_{i=1}^{N_{test}}$) from the training set for every pending solving function. We denote the similarity values between $a^{test}_{i}$ and $a^{test}_{k_{i}}$ as $s^{test}_{i}$. (2) Calculate the maximum value $s^{test}_{max}$ of $\{s^{test}_{i}\}_{i=1}^{N}$. (3) For every training trajectory $a^{train}_{i}$, calculate its similarity with other training trajectories. Denote the index of $r$-th similar trajectory with $a^{train}_{i}$ as $k^{r}_{i}$, and the similar score between $a^{train}_{i}$ and $a^{train}_{k^{r}_{i}}$ as $s^{r}_{i}$. (4) The $K$ is calculated as the minimum $r$ satisfying $s^{test}_{max}\leq s_{i}^{r}$ for any $i\in[1,N^{train}]$.
C.2 Variant Resolution Auxiliary Trajectory Integration
Sometimes the auxiliary function $a_{k_{i}}$ has a different resolution. For example, when we need to predict high-resolution physical trajectory using low-resolution trained models, the auxiliary samples $a_{k_{i}}$ in the training set are more coarse-grained than the queried function $a^{H}$. In this scenario, we could interpolate the low-resolution auxiliary sample to a fine-grained grid, then concatenate the interpolated high-resolution grid with the queried function $a^{H}_{i}$.
$a_{input}=\mathcal{F}^{inter}(a_{k_{i}})\oplus a^{H}.$ (13)
In this work, we implement the interpolation method $\mathcal{F}^{inter}$ based on the Fourier Transform. Specifically, we first transform the coarse-grained input $u_{k_{i}}$ into Fourier Space, fill zeros in the high-frequency region to target resolution, and finally transform it back to spatial space. Our experimental results show the trained residual neural operator could generalize well in this interpolation scenario.
C.3 Variant Resolution Trajectory Retrieval
For the high resolution inference scenario, we could first downsample the high-resolution queried function $a^{H}$ to low resolution $a^{H}_{low}$, then calculate the similarity score:
$\mathcal{F}^{ret}(a^{H},\mathcal{T})$ $=\text{argmax}_{j}\,\text{Sim}(a^{H}_{low},a_{j}),$ (14)
$a^{H}_{low}$ $=\mathcal{F}^{low}(a^{H}),$
where $\mathcal{F}^{low}$ is down sampling function. In this work, we implement $\mathcal{F}^{low}$ by erasing additional information in high-frequency regions based on the Fourier Transform.
C.4 Complexity Analysis of Physical Trajectory Retrieval
The computational cost for retrieval is negligible compared to neural network inference. Here is a detailed analysis of the computational cost for physical trajectory retrieval.
The computational expensive is mainly related to two factors: the training set size (noted as $N_{train}$) and the number of physical field points (ie. the resolution of physical field) (noted as $N_{field}$). The computing process consists of the following two steps:
  • Step 1: Similarity Score Calculation. For every sample in the retrieval set, the similarity metric function (eg. Cosine Similarity) could be calculated with linear time complexity $O(N_{field})$. Thus the time complexity of calculating similarity scores for all retrieved samples is $O(N_{field}\cdot N_{train})$.
  • Step 2: Ranking based on similarity scores. We use the simple sorting algorithm (Quicksort) to rank all samples based on their similarity score. The time complexity is $O(N_{train}\cdot\log N_{train})$. The overall time complexity for retrieval is $O(N_{field}\cdot N_{train}+N_{train}\cdot\log N_{train})$.
Commonly, the numerical ranges for $N_{field}$ and $N_{train}$ are approximately between $10^{3}$ to $10^{4}$, and $10^{2}$ to $10^{4}$, respectively. Therefore, the real computational amount is about $O(10^{8})$, this could be calculated on most computers. In our implementation, the average retrieval time on Darcy Flow ($85\times 85$ resolution) is 2.67e-3 to 1.84e-2 seconds for different retrieval set sizes (100 to 900).
Appendix D Visualization Analysis Detail
D.1 Similarity between $u(x)$ Visulization Detail
We visualize the correlation between $u(x)$ normalized distance and $a(x)$ similarity rank (shown in Figure 4) as following steps: Firstly, taking the training set $\mathcal{T}$ as retrieval set, we sample auxiliary trajectory $(a_{k_{i,r}},u_{k_{i,r}})$ with $r$-th top similarity score for every test sample $(a^{test}_{i},u^{test}_{i})$ in the test set. Here $r$ represents the similarity rank (higher $r$ indicates lower similarity), and we take gradually increasing values from $1$ to $80$ for $r$. Next, we calculate the normalized distance between $u^{test}_{i}$ with its auxiliary sample $u_{k_{i,r}}$. The normalized distance is defined the same as the relative error Equation 14.
D.2 Label Distribution Visualization Detail
We visualize the label distribution (as shown in  Figure 5) by reducing the high-dimension function fields to 2 dimensions utilizing Principle Component Analysis. Specifically, for both direct operator mapping and residual operator mapping, we first optimize a PCA model on the training labels (100 labels) and then calculate the dimension-reduced labels for every trajectory in the training and testing set. Finally, we visualize every two-dimension label point on the graph as well as the ellipses representing the stand deviation and the range along each dimension.

### v2

@@
-DeltaPhi: Learning Physical Trajectory Residual for PDE Solving
+DeltaPhi: Physical States Residual Learning for Neural Operators in Data-Limited PDE Solving
 Abstract
-Although neural operator networks theoretically approximate any operator mapping, the limited generalization capability prevents them from learning correct physical dynamics when potential data biases exist, particularly in the practical PDE solving scenario where the available data amount is restricted or the resolution is extremely low.
-To address this issue, we propose and formulate the Physical Trajectory Residual Learning (DeltaPhi), which learns to predict the physical residuals between the pending solved trajectory and a known similar auxiliary trajectory.
-First, we transform the direct operator mapping between input-output function fields in original training data to residual operator mapping between input function pairs and output function residuals.
-Next, we learn the surrogate model for the residual operator mapping based on existing neural operator networks.
-Additionally, we design helpful customized auxiliary inputs for efficient optimization.
-Through extensive experiments, we conclude that, compared to direct learning, physical residual learning is preferred for PDE solving.
+The limited availability of high-quality training data poses a major obstacle in data-driven PDE solving, where expensive data collection and resolution constraints severely impact the ability of neural operator networks to learn and generalize the underlying physical system.
+To address this challenge, we propose DeltaPhi, a novel learning framework that transforms the PDE solving task from learning direct input-output mappings to learning the residuals between similar physical states, a fundamentally different approach to neural operator learning.
+This reformulation provides implicit data augmentation by exploiting the inherent stability of physical systems where closer initial states lead to closer evolution trajectories.
+DeltaPhi is architecture-agnostic and can be seamlessly integrated with existing neural operators to enhance their performance.
+Extensive experiments demonstrate consistent and significant improvements across diverse physical systems including regular and irregular domains, different neural architectures, multiple training data amount, and cross-resolution scenarios, confirming its effectiveness as a general enhancement for neural operators in data-limited PDE solving.
 1 Introduction
-Due to lacking analytical solutions, solving complex Partial Differential Equations (PDEs), e.g.
-Navier-Stokes, relies on numerical simulation based on ultra-fine grid division, which consumes expensive computational resources and time.
-To resolve this, machine learning based PDE solving [1, 2, 3, 4, 5] has sparked extensive research interest in recent years.
-Notably, the operator learning [6, 2, 1] directly learns the mapping between the input-output function fields, forming a surrogate model for a given parametric PDE.
-Then the solution for any input functions could be quickly obtained via neural network forward computation.
-Despite current operator learning networks [2, 3] theoretically being able to approximate any operator mapping, their generalization capabilities remain limited.
-First, the networks are prone to overfitting to training resolution while suffering performance degradation on other resolutions [7], especially when the training resolution is extremely low.
-Second, when available training data is limited, the network may struggle to accurately model the correct physical dynamics due to potential biases in the data.
-Third, data scarcity and bias are common occurrences in practical PDE solving [8, 9].
-Therefore, further investigation is required to enhance the neural operator generalization ability.
-In this work, for generalized PDE solving, we introduce the Physical Trajectory Residual Learning (DeltaPhi), which reformulates the direct operator mapping between the input-output function fields as the mapping between input function pairs and output function residuals (residual operator mapping).
-The residual operator mapping could be learned with any existing neural operators (such as FNO [2], NORM [10], FFNO [11], etc.
-), thereby obtaining corresponding residual neural operator.
-First, we reconstruct the training set from data pairs of direct operator mapping to data pairs of residual operator mapping, via assigning an auxiliary trajectory to every training data.
-The auxiliary trajectory is a different trajectory sampled from the training set based on similar trajectory retrieval.
-Next, we optimize the neural operator network with reconstructed training data.
-By concatenating the pending solved input function and auxiliary trajectory together as input of the network, and adding the auxiliary trajectory solution to the final output of the network via residual connection, we could optimize the network with existing training frameworks.
-All other setups (including the loss function, trainable weights amount, etc.
-) remain consistent with the base operator networks.
-Compared to previous direct operator learning, the reconstructed residual training labels could be customized with enhanced diversity through selecting appropriate auxiliary samples.
-Consequently, the restricted diversity and potential distribution bias due to limited training data amount could be mitigated, thereby obtaining more generalized surrogate models.
-Residual learning methodology has been extensively studied in the machine learning community [12, 13, 14, 15].
-Previous works present various benefits derived from residual learning, such as enhancing generalization capability, handling complex regression mapping, etc.
-In this work, we introduce this concept into the neural operator learning task, a more complex regression task involving learning the mapping between high-dimension (tens of thousands) fields.
-The residuals between different but similar physical trajectories are learned, and utilized to solve PDEs.
-Through the visualization analysis of label distribution, we statistically validate the favorable generalization property of learning the residual operator mapping.
-The contributions are summarized as follows:
-  • We propose and formulate a novel framework for data-driven PDE solving in restricted data (limited amount, low resolution, biased distribution) scenarios, i.e.
-the Physical Trajectory Residual Learning (DeltaPhi).
-  • We statistically analysis the advantage of residual operator learning and make instantiations based on similar trajectory retrieval and existing neural operator networks.
-  • We validate the superiority of residual operator learning through extensive experiments across different PDE problems, different physical domains (both regular and irregular problems), different base models, different training set sizes, and different training resolutions.
-Additionally, detailed ablation studies are conduct.
-2 Related Work
-Neural operator learning.
-Fourier Neural Operator (FNO) [2] utilizes the Fourier Transform based integral operation to efficiently learn the physical dynamics using limited training data.
-Factorized Fourier Neural Operator (FFNO) [11] factorize the Fourier Transform along each dimension, reducing the number of network weights.
-Clifford Fourier Neural Operator (CFNO) [16] employs the Clifford Algebra in the network architecture, incorporating geometry prior between multi physical fields.
-The various neural operators could be simply integrated with the proposed Physical Trajectory Residual Learning.
-Other works investigate architecture improvements for different needs, including chaotic systems modeling [17], physical-informed instance-wise finetuning [7], accelerating computation [18, 19, 20], spherical fields processing [21], irregular fields learning [22, 23], non-periodic boundary fields modeling [24], large-scale pretraining [25, 26], etc.
+Due to the lack of analytical solutions, solving complex Partial Differential Equations (PDEs) e.g.
+Navier-Stokes traditionally relies on numerical simulations based on ultra-fine grid division, which consumes expensive computational resources and time.
+This computational bottleneck has motivated the development of machine learning based PDE solving approaches (kovachki2021neural, ; li2020fourier, ; lu2019deeponet, ; raissi2019physics, ; long2018pde, ).
+In particular, neural operator learning (kovachki2021neural, ; li2020fourier, ; li2020neural, ) has emerged as a promising direction by directly learning the mapping between input-output function fields, enabling efficient solution prediction through neural network forward computation.
+A critical challenge in neural operators is the severe scarcity of high-quality training data (bonnet2022airfrans, ; benitez2024out, ), which manifests in multiple aspects: (1) collecting comprehensive training data often require expensive physical experiments or high-resolution numerical simulations; (2) computational and storage constraints may restrict the resolution of available data; (3) the collected data frequently exhibits distribution bias due to physical or experimental limitations.
+These data limitations fundamentally constrain the generalization capability of current neural operators.
+To address these challenges, we propose DeltaPhi, a novel framework that helps neural operators learn from scarce data.
+The core innovation of DeltaPhi lies in reformulating the PDE solving task from learning direct input-output mappings to learning the residuals between similar physical states.
+This reformulation enables implicit data augmentation by leveraging a key property of physical systems: states with similar initial conditions tend to follow similar solutions.
+By learning residuals between pairs of similar trajectories, DeltaPhi effectively expands the diversity of training samples without requiring additional data collection.
+This is fundamentally different from existing approaches that attempt to learn the straight mapping for each sample independently.
+The proposed framework presents two advantages: (1) Enhanced Training Sample Diversity: The framework enables enhanced diversity in training labels through random auxiliary solution sampling, effectively mitigating distribution bias issues.
+This training distribution augmentation preserves physical validity while expanding the learned solution space.
+(2) Architecture Flexibility: DeltaPhi is architecture-agnostic and can seamlessly enhance existing neural operators (li2020fourier, ; chen2023learning, ; tran2021factorized, ) through a unified residual learning mechanism, making it applicable across different PDE solving scenarios.
+We conduct extensive experiments to validate the effectiveness of physical states residual learning across various settings: (1) diverse physical systems including both regular domains (Darcy Flow, Navier-Stokes) and irregular domains (Heat Transfer, Blood Flow); (2) different neural architectures including spectral-based and attention-based operators; (3) varied numbers of training samples to test robustness under data scarcity; and (4) cross-resolution scenarios to evaluate generalization capability.
+The results consistently demonstrate that DeltaPhi significantly improves the performance of base models, particularly in data-limited settings.
+Overall, this work introduces a promising framework for enhancing neural operators in practical PDE solving applications where high-quality training data is scarce or expensive to obtain.
+The physical states residual learning framework could potentially benefit other high-dimensional regression problems in scientific machine learning, especially when dealing with physically governed systems that exhibit stability properties.
+2 Background and Related Work
+2.1 Direct Neural Operator Learning
+Neural operator learning aims to learn mappings between infinite-dimensional function spaces for solving parametric PDEs (kovachki2021neural, ).
+The operator mapping $\mathcal{G}:\mathcal{A}\longrightarrow\mathcal{U}$ maps between Banach spaces $\mathcal{A}$ and $\mathcal{U}$.
+For steady-state problems like Darcy Flow, the input function $a(x)\in\mathcal{A}$ typically represents coefficients while $u(x)\in\mathcal{U}$ is the solution.
+For time-series problems like Navier-Stokes, $a(x)$ commonly represents prior states, and $u(x)$ represents future states (li2020fourier, ).
+Neural operators are constructed using various network architectures, with two prominent families being kernel-based operators and DeepONet-style operators.
+KernelIntegral-form architectures, such as Graph Neural Operators (GNO) (kovachki2021neural, ), are often based on iterative kernel integration.
+Deep Operator Networks (DeepONet) (lu2019deeponet, ) instead use separate branch and trunk networks to approximate the operator.
+These two general forms can be represented as:
+KernelIntegral-form: $\quad\mathcal{G}_{\theta}=Q\circ\sigma(W_{l}+\mathcal{K}_{l})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1})\circ P$ (1)
+DeepOnet-form: $\quad\mathcal{G}_{\theta}(a)(y)=\sum_{k=1}^{p}B_{k}(a)T_{k}(y)$
+In the KernelIntegral-form, $P$ and $Q$ are projection layers, $\mathcal{K}_{i}$ are learnable kernel integral operators, $W_{i}$ are local linear operators, and $\sigma$ is a nonlinear activation.
+In the DeepOnet-form, $B_{k}(a)$ represents the branch network outputs encoding the input function $a$, and $T_{k}(y)$ represents the trunk network outputs encoding the output coordinates $y$.
+It is noted that the framework proposed in this work is applicable to both forms of architectures.
+The components of these architectures are differently instantiated in previous works.
+Fourier Neural Operator (FNO) (li2020fourier, ) utilizes the Fourier Transform based integral operation to efficiently learn the physical dynamics using limited training data.
+Factorized Fourier Neural Operator (FFNO) (tran2021factorized, ) factorize the Fourier Transform along each dimension, reducing the number of network weights.
+Clifford Fourier Neural Operator (CFNO) (brandstetter2022clifford, ) employs the Clifford Algebra in the network architecture, incorporating geometry prior between multi-physical fields.
+The various neural operators could be simply integrated with the proposed Physical States Residual Learning.
+Other works investigate architecture improvements for different needs, including chaotic systems modeling (li2021learning, ), physical-informed instance-wise finetuning (li2021physics, ), accelerating computation (poli2022transform, ; white2023speeding, ; rahman2022u, ; xiao2025quantum, ; DBLP:journals/jcphy/GaoW23, ), spherical fields processing (bonev2023spherical, ), irregular fields learning (li2022fourier, ; hao2023gnot, ), non-periodic boundary fields modeling (liu2023spfno, ), large-scale pretraining (grady2023model, ; subramanian2023towards, ), etc.
 In addition, some works explore other network backbones, e.g.
-improved frequency-spatial domain transformation [27, 28, 29, 30], convolutions [31], graph neural network [32, 6, 33], attention mechanism [34, 23, 35], and diffusion models [36].
-Incorporating residual learning with these designs is worth exploring in the future.
-Residual learning in previous works.
-Some works learn residual between different data samples.
-Similar to ResMem [12], we also utilize the training set as auxiliary memory during inference, enjoying its generalization enhancement property.
+improved frequency-spatial domain transformation (yueholistic, ; gupta2021multiwavelet, ; zhao2022incremental, ; cao2023lno, ; tripura2023wavelet, ; wang2024latent, ; gaodynamic, ; gao2025discretization, ), convolutions (raonic2023convolutional, ), graph neural network (brandstetter2022message, ; li2020neural, ; li2020multipole, ), attention mechanism (cao2021choose, ; hao2023gnot, ; prasthofer2022variable, ; wu2024transolver, ), and diffusion models (lim2023score, ; zhou2024unisolver, ; shan2025red, ; shan2024pird, ).
+Despite theoretical approximation capabilities, practical limitations persist in generalization across resolutions (li2021physics, ), handling data scarcity (vinuesa2022emerging, ), and managing biased training distributions (bonnet2022airfrans, ).
+Our work addresses these challenges through residual learning between physical trajectories.
+2.2 Residual Learning in Previous Works
+Some works (yang2023resmem, ; xu2023sequential, ; zhang2018residual, ; freund1995desicion, ) learn residual between different data samples.
+Similar to ResMem (yang2023resmem, ), we also utilize the training set as auxiliary memory during inference, enjoying its generalization enhancement property.
 The difference lies that (1) we memorize the auxiliary labels and learn the residuals, while ResMem memorizes the residuals and learns the labels, (2) our framework is end-to-end trained while ResMem is trained in separate stages.
-Similar external memory augmented strategies are popular in the Language Modeling community [37, 38].
-Other works learn residuals between low resolution and high resolution representation for image super-resolution [39] and deep learning based CFD simulation [39].
-Similar residual learning is investigated in time series prediction task  [40, 13].
+Similar external memory augmented strategies are popular in the Language Modeling community (wang2022training, ; guu2020retrieval, ).
+Other works learn residuals between low resolution and high resolution representation for image super-resolution (jessica2023finite, ) and deep learning based CFD simulation (jessica2023finite, ).
+Similar residual learning is investigated in time series prediction task  (angelopoulos2023conformal, ; xu2023sequential, ).
 The primary distinction between these works to ours is that we do not utilize the corresponding low-resolution solutions or previous time-step solutions, but instead, learn the residuals between different trajectories.
+Additionally, (cheng2024reference, ) studies learning residuals between PDE solutions with similar geometric domains.
+In contrast, our work focuses on a fundamentally different problem of learning residuals between similar physical states, enabling broader applicability.
+We establish this based on physical system stability and develop implicit data augmentation through similarity-based sampling during training, effectively addressing distribution bias and overfitting issues in data-limited scenarios.
 3 Methodology
-3.1 Preliminary
-3.1.1 Direct learning formulation
-Direct operator mapping. Previous works [2, 11, 23] formulate the parametric PDE numerical solution as an operator mapping between two infinite dimensional function spaces:
-$\mathcal{G}:\mathcal{A}\longrightarrow\mathcal{U},$ (1)
-where $\mathcal{A}$ and $\mathcal{U}$ are both banach spaces of functions.
-For steady-state solving problems, e.g.
-Darcy Flow, $a(x)\in\mathcal{A}$ represents the diffusion coefficient function, and $u(x)\in\mathcal{U}$ is the corresponding solution function.
-For time-series forecasting problems, e.g.
-Navier-Stokes, $a(x)\in\mathcal{A}$ typically denotes the state function of the preceding few steps, while $u(x)\in\mathcal{U}$ represents the function in the subsequent few steps.
-The learning task is to study a parameterized surrogate model $\mathcal{G}_{\theta}$ for $\mathcal{G}$.
-Domain discretization.
-Following [2, 41], we numerically process the function $a(x)$ and $u(x)$ by taking the discretization of the continuous domain.
-For regular domain problems, we uniformly discretize the regular field, using standard rectangular grids to capture the physical states.
-For irregular domain problems, we take the discrete mesh representation of the field.
-Optimization problem.
-Consider the training set $\mathcal{T}$ consisting of $N$ i.i.d.
-pairs $\{a_{i},u_{i}\}_{i=1}^{N}$, where $a_{i}\in\mathcal{A}$, $u_{i}\in\mathcal{U}$, the objective is to find $\theta$ that minimizes the following cost function:
-$\ell=\sum_{i=1}^{N}\mathcal{L}(\mathcal{G}_{\theta}(a_{i}),u_{i}),$ (2)
-where $\mathcal{L}$ is a cost function that measures the difference between two functions.
-During inference, given the input function $a\in\mathcal{A}$, we obtain the corresponding solution $u=\mathcal{G}_{\theta}(a)$.
-3.1.2 Neural Operator Network
-Prior works [2, 34, 11] have extensively explored neural network architectures for learning the operator mapping $\mathcal{G}_{\theta}$ between function spaces.
-Among these, the Fourier Neural Operator (FNO)  [2, 1] has been the most commonly applied model across various physical systems.
-Neural operators are defined as multi-layer networks formulated as follows:
-$\mathcal{G}_{\theta}=Q\circ\sigma(W_{l}+\mathcal{K}_{l})\circ\cdots\circ\sigma(W_{i}+\mathcal{K}_{i})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1})\circ P,$ (3)
-where $l$ represents the number of layers in the network.
-The projector $P$ encodes the input function $a_{i}$ into an implicit embedding $v_{0}$.
-The operators $W_{i}$ and $\mathcal{K}_{i}$ at the $i$-th layer transform the input function $v_{i-1}$ into $v_{i}$.
-$\sigma$ is an element-wise non-linear operation introduced for learning non-linear mapping between function spaces.
-$Q$ is the output projector which decodes the last layer’s output $v_{l}(x)$ into the targeted function $u_{i}(x)$.
-The components of $\mathcal{G}_{\theta}$ could be differently instantiated across various models.
-Typically, $P$ and $Q$ are implemented as fully connected layers.
-In Fourier Neural Operator [2], $\mathcal{K}_{i}$ is instantiated as an integral operator based on Fourier Transform, and $W_{i}$ is a linear fully connected layer.
-In NORM [10], $\mathcal{K}_{i}$ is implemented based on the Laplacian Transform for processing irregular domains.
-In addition, $\mathcal{K}_{i}$ can also be a convolution operation, adopting an architecture similar to Resnet [42].
-Although direct operator learning networks can theoretically approximate any operator mapping, they encounter limited generalization capabilities in practical applications.
-First, the networks tend to overfit on trained resolution, and suffer performance drops on high resolution inference.
-Second, the model is vulnerable to potential bias in training data (especially when the data amount is limited), thus failing to capture the intrinsic physical trajectory dynamics.
-3.2 Physical Trajectory Residual Learning
-Figure 1: The overall architecture of Physical Trajectory Residual Learning.
-Given an input trajectory $a_{i}$, we first sample a similar auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$ from the training trajectory set $\mathcal{T}$.
-Subsequently, the trajectory $a_{i}$ and $(a_{k_{i}},u_{k_{i}})$ are concatenated and fed into the the neural operator $\mathcal{G}_{\theta}^{\Delta}$, producing the predicted trajectory residual.
-Finally, the predicted solution $\hat{u}_{i}$ is obtained by adding the predicted trajectory residual with the auxiliary solution $u_{k_{i}}$.
-3.2.1 Formulation
-Different from the aforementioned direct operator learning, we propose to learn the residual operator mapping that captures the differences between pairs of physical trajectories:
+3.1 Preliminary: Stability of Physical Systems
+Stability of PDEs. For a well-posed PDE system with solution operator $\mathcal{G}$, stability means:
+$\|\mathcal{G}(a_{1})-\mathcal{G}(a_{2})\|_{\mathcal{U}}\leq C\|a_{1}-a_{2}\|_{\mathcal{A}},$ (2)
+where $a_{1},a_{2}$ are input functions and $C$ is a positive constant.
+This property, also known as Lipschitz continuity of the solution operator, is one of Hadamard’s criteria for well-posedness of PDEs.
+This stability property is well-established in various PDEs: elliptic PDEs satisfy stability under appropriate boundary conditions; parabolic PDEs exhibit stability through their dissipative mechanisms; and even certain chaotic systems maintain stability for suitable time intervals.
+It’s noted that this stability condition is not an additional assumption imposed by our method, but rather a fundamental prerequisite for any data-driven operator learning approach to succeed (lu2019deeponet, ; li2020fourier, ).
+If small changes in the input function could lead to arbitrarily large, discontinuous changes in the solution, learning a generalizable mapping would be extremely challenging for neural networks.
+Foundation of Residual Learning.
+Stability of PDEs establishes a relationship where the magnitude range of difference between output functions systematically changes with respect to the difference between input functions.
+This relationship allows us to explicitly control the diversity of output function residuals by selecting input function pairs with varying similarity ranges.
+This establishes the foundation of residual neural operator learning and enables the effectiveness of implicit data augmentation, introduced in Section˜3.5.
+3.2 Residual Operator Mapping: A Unified Formulation
+Figure 1: The overall architecture of Physical States Residual Learning.
+Given an input function $a_{i}$, we first sample a similar auxiliary sample $(a_{k_{i}},u_{k_{i}})$ from the training sample set $\mathcal{T}$.
+Subsequently, $a_{i}$ and $(a_{k_{i}},u_{k_{i}})$ are concatenated and fed into the the neural operator $\mathcal{G}_{\theta}^{\Delta}$, producing the predicted states residual.
+Finally, the predicted solution $\hat{u}_{i}$ is obtained by adding the predicted states residual with the auxiliary solution $u_{k_{i}}$.
+Residual Operator Mapping. Based on the stability of physical systems, we formulate the residual operator mapping that captures the differences between pairs of physical trajectories:
 $\mathcal{G}^{\Delta}$ $:\mathcal{A}^{2}\longrightarrow\Delta\,\mathcal{U},$ (3)
-$\mathcal{A}^{2}$ :$:\{(a_{i},a_{j})\mid a_{i}\in\mathcal{A},a_{j}\in\mathcal{A}\},$
-$\Delta\,\mathcal{U}$ :$:\{u_{i}-u_{j}\mid u_{i}\in\mathcal{U},u_{j}\in\mathcal{U}\},$
+$\mathcal{A}^{2}$ $:\{(a_{i},a_{j})\mid a_{i}\in\mathcal{A},a_{j}\in\mathcal{A}\},$
+$\Delta\,\mathcal{U}$ $:\{u_{i}-u_{j}\mid u_{i}\in\mathcal{U},u_{j}\in\mathcal{U}\},$
 where $\mathcal{G}^{\Delta}$ represents the residual mapping operator.
@@
 $a_{*}$ and $u_{*}$ are functions in the space of $\mathcal{A}$ and $\mathcal{U}$, respectively.
-The residual mapping operator $\mathcal{G}^{\Delta}$ maps the input functions pair $(a_{i},a_{j})$ into the corresponding function residual $u_{i}-u_{j}$, i.e.
-$u_{i}-u_{j}=\mathcal{G}^{\Delta}(a_{i},a_{j})$.
-This approach leverages the concept of residuals to simplify the learning task.
-Instead of learning the direct operator mapping from $\mathcal{A}$ to $\mathcal{U}$, we focus on learning the differences between output functions.
-The transformed training label distribution, i.e.
-the function residual distribution, could have sufficient diversity by selecting appropriate auxiliary samples, thereby alleviating the training data bias and better handling the unknown function to be solved.
-We detail the explanation about this in Section 3.2.5.
-Both direct operator mapping and residual operator mapping are mappings between Banach spaces of functions.
-Therefore, we could employ the off-the-shelf neural operator networks [2, 11] as the base network backbone.
-Neural operator networks are designed to handle mapping between function spaces and are known for their universal approximation capabilities, making them well-suited for residual operator mapping.
-3.2.2 Auxiliary Trajectory Sampling
-Given an input function (or named as input trajectory) $a_{i}$, we need an auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$ to learn the residual neural operator $\mathcal{G}^{\Delta}$ and infer the corresponding output $u_{i}$ of $a_{i}$.
-The auxiliary trajectory can be sampled from the existing data, allowing us to leverage prior knowledge to improve the generalization capabilities of the model.
-By selecting an appropriate auxiliary trajectory, we can simplify the computation and enhance the robustness of the residual operator mapping.
-Given the training set $\mathcal{T}=\{a_{i},u_{i}\}_{i=1}^{N}$, each data pair $(a_{i},u_{i})$ is associated with an auxiliary trajectory $(a_{k_{i}},u_{k_{i}})$, sampled from $\mathcal{T}$.
-The index $k_{i}$ of the auxiliary trajectory is defined by:
-$k_{i}=\mathcal{F}^{ret}(a_{i},\mathcal{T}),\,k_{i}\in\{1,2,...,N\},\,k_{i}\neq i,$ (5)
-where $\mathcal{F}^{ret}$ is the sampling function.
-A straightforward implementation of $\mathcal{F}^{ret}$ could be random sampling from $\{1,2,...,i-1,i+1,...N\}$.
-Using the calculated $k_{i}$, we transform the origin training set $\mathcal{T}$ to a residual training set $\mathcal{T}^{res}=\{(a_{i},a_{k_{i}}),u_{i}-u_{k_{i}}\}_{i=1}^{N}$.
-Below we present our implementation of $\mathcal{F}^{ret}$ in detail.
-Training stage.
-During training, give an input function $a_{i}$, we randomly sample a trajectory $a_{j}$ from the set of trajectories with top-k similarity scores to $a_{i}$.
-The sampling function $\mathcal{F}^{ret}$ is defined as:
-$\mathcal{F}^{ret}(a_{i},\mathcal{T})$ $=\texttt{RandSample}(\texttt{argtopK}_{j,j\neq i}\,\texttt{Sim}(a_{i},a_{j})),$ (6)
-$\texttt{Sim}(a_{i},a_{j})$ $=1-{a_{i}\cdot a_{j}\over\|a_{i}\|\|a_{j}\|},$
+Solving PDEs by Residual Operator $\mathcal{G}^{\Delta}$.
+In PDE solving, given an input function $a_{i}$, we first retrieve an auxiliary sample $(a_{k_{i}},u_{k_{i}})$ from the set of physical samples with known solution e.g.
+training set.
+Then the residual neural operator $\mathcal{G}^{\Delta}$ predicts the solution residual between $u_{i}$ and $u_{k_{i}}$.
+The final solution $u_{i}$ could be obtained by adding this predicted residual to $u_{k_{i}}$:
+$u_{i}=\mathcal{G}^{\Delta}(a_{i},a_{k_{i}})+u_{k_{i}}.$ (4)
+We introduce how to obtain auxiliary samples in Section˜3.3, and the specific implementation of residual neural operators in Section˜3.4.
+3.3 Auxiliary Sample Retriever $\mathcal{F}^{\text{ret}}$
+Given an input function $a_{i}$ and the training set $\mathcal{T}=\{a_{i},u_{i}\}_{i=1}^{N}$, we could retrieve an auxiliary sample $(a_{k_{i}},u_{k_{i}})$ from $\mathcal{T}$, formulated as follows:
+$k_{i}=\mathcal{F}^{\text{ret}}(a_{i},\mathcal{T}),\,k_{i}\in\{1,2,...,N\},\,k_{i}\neq i,$ (5)
+where $\mathcal{F}^{\text{ret}}$ is the sampling function. Below we present our implementation of $\mathcal{F}^{\text{ret}}$ in detail.
+Training stage. During training, given an input function $a_{i}$, we randomly retrieve a sample $a_{j}$ from training set with top-k similarity scores to $a_{i}$:
+$\mathcal{F}^{\text{ret}}(a_{i},\mathcal{T})$ $=\texttt{RandSample}(\texttt{argtopK}_{j,j\neq i}\,\texttt{Sim}(a_{i},a_{j})),$ (6)
+$\texttt{Sim}(a_{i},a_{j})={a_{i}\cdot a_{j}\over\|a_{i}\|\|a_{j}\|},$
 where $\texttt{RandSample}(\cdot)$ represents randomly sampling one element from the given set.
@@
 For a fair comparison, except for specific statements, we set $K=20$ for all experiments in this work.
-Inference stage. During inference, given the input function $a^{test}$, we select the most similar auxiliary trajectory from the training set:
-$\mathcal{F}^{ret}(a^{test},\mathcal{T})=\texttt{argmax}_{j}\,\texttt{Sim}(a^{test},a_{j}).$ (7)
-We empirically found the model performance is robust to the choice of auxiliary trajectories.
-In Section A.3, we experiment with randomly sampled auxiliary trajectories from the top ten similarity scores and find the standard deviation across five repeating evaluations is quite small.
-This indicates that the residual neural operator effectively learns the residual operator mapping for the same $a_{i}$ with varied auxiliary trajectories.
-We further present the trajectory sampling for varied resolution fields in Section C.3.
-Section C.4 provides the cost analysis of the trajectory sampling.
-3.2.3 Auxiliary Trajectory Integration
-As aforementioned, the residual operator $\mathcal{G}^{\Delta}$ requires additional auxiliary function $a_{k_{i}}$ as input.
+Inference stage. During inference, given the input function $a^{test}$, we select the most similar auxiliary sample from the training set:
+$\mathcal{F}^{\text{ret}}(a^{test},\mathcal{T})=\texttt{argmax}_{j}\,\texttt{Sim}(a^{test},a_{j}).$ (7)
+While we take the most similar sample for inference, Section˜A.4 shows that the model is robust to different auxiliary sample choices.
+When randomly selecting from the top 10 similar samples, repeated testing shows minimal variation - a standard deviation of just 5.41e-5 across five runs.
+This indicates that neural operators effectively learn to handle varied auxiliary samples for the same input.
+Cross-Resolution Sample Retrieval.
+For cross-resolution scenarios where auxiliary samples have different resolutions from the input function, we employ Fourier Transform-based alignment.
+Specifically, during inference with high-resolution input function $a^{H}$, we downsample $a^{H}$ to match training resolution via truncating high-frequency components for similarity calculation:
+$\mathcal{F}^{\text{ret}}_{\text{aligned}}(a^{H},\mathcal{T})=\mathcal{F}^{\text{ret}}(\text{DownSampling}(a^{H}),a_{j}),$ (8)
+where DownSampling is implemented through frequency-domain truncation:
+$\text{DownSampling}(a^{H})=\mathcal{F}_{\text{fourier}}^{-1}(\text{Truncate}(\mathcal{F}_{\text{fourier}}(a^{H}))).$ (9)
+Here $\mathcal{F}_{\text{fourier}}$ and $\mathcal{F}_{\text{fourier}}^{-1}$ denote the Fourier transform and its inverse, respectively.
+The Truncate operation removes high-frequency components in Fourier space to match the target resolution.
+This Fourier-based downsampling preserves the low-frequency structures of the physical field while ensuring consistent resolution for similarity computation.
+This resolution alignment enables effective utilization of auxiliary samples across different resolutions.
+Retrieval Cost Analysis.
+The computational overhead of our retrieval process is minimal compared to neural network inference.
+The total complexity consists of similarity score calculation ($O(N_{field}\cdot N_{train})$) and ranking ($O(N_{train}\cdot\log N_{train})$), where $N_{field}$ is the number of field points and $N_{train}$ is the training set size.
+Our comprehensive experiments on Darcy Flow demonstrate the high efficiency: GPU-based retrieval only takes 0.2-0.3ms across different training set sizes (100-900), and the entire residual learning pipeline, including both retrieval and data preparation, adds merely 0.5ms to the inference time.
+This negligible overhead makes our method highly practical for real-world applications.
+Detailed complexity analysis and timing experiments are provided in Section˜C.2.
+{algorithm}
+[…]
+3.4 Architecture of Residual Neural Operators $\mathcal{G}_{\theta}^{\Delta}$
+Residual Neural Operator Backbone.
+Benefiting from the universal approximation capability of neural operator networks (li2020fourier, ; tran2021factorized, ) for direct operator learning, they could be quickly employed as residual neural operators $\mathcal{G}_{\theta}^{\Delta}$ for learning residual operator mapping $\mathcal{G}^{\Delta}$:
+$\mathcal{G}_{\theta}^{\Delta}=Q\circ\sigma(W_{l}+\mathcal{K}_{l})\circ\cdots\circ\sigma(W_{1}+\mathcal{K}_{1})\circ P,$ (10)
+where $P$ and $Q$ are projection layers, $\mathcal{K}_{i}$ and $W_{i}$ are learnable operators, and $\sigma$ is a nonlinear activation.
+Next, we introduce how to modify existing neural operators for learning residual mappings.
+Concatenation of Auxiliary Samples.
+As aforementioned, the residual operator $\mathcal{G}^{\Delta}_{\theta}$ requires additional auxiliary function $a_{k_{i}}$ as input.
 In specific implementation, $a_{i}$ and $a_{k_{i}}$ could be straightly concatenated along their channel dimension:
 $a_{input}=a_{i}\oplus a_{k_{i}},$ (11)
-where $\oplus$ represents the concatenation operation of two vectors. Then the concatenated vector $a_{input}$ is feeded into the neural operator network:
+where $\oplus$ represents the concatenation operation of two vectors. Then the concatenated vector $a_{input}$ is feed into the neural operator network:
 $v_{0}=P(a_{input}),$ (12)
-where $P$ is the fully connected encoder in the operator network (Equation 3), and $v_{0}$ represents the first latent function in the hidden space.
-Integration of auxiliary trajectory with variant resolution.
-In cases where the auxiliary input function $a_{k_{i}}$ has a different resolution from the input function $a_{i}$, we address this discrepancy via field interpolation techniques.
-We present the details in Section C.2.
-Integration of auxiliary trajectory with additional related information.
+where $P$ is the fully connected encoder in the operator network (Equation˜10), and $v_{0}$ represents the first latent function in the hidden space.
+Concatenation of Additional Information.
 The residual operator mapping can be challenging to learn due to the variability of output residuals for different auxiliary samples.
-To mitigate such difficulty, we introduce additional information related to $a_{k_{i}}$ as inputs for residual neural operators:
-  • For time-series PDE solving, e.g. Navier-Stokes, we only input the last few time steps of $a_{k_{i}}$.
-  • We use the corresponding $u_{k_{i}}(x)$ of $a_{k_{i}}(x)$ as the input.
-  • We utilize calculated relation metrics, e.g. similarity between $a_{i}$ and $a_{k_{i}}$, as inputs.
-Our experimental results (Section A.2) indicate that this customized input strategy indeed improves the performance of residual neural operators.
-3.2.4 Residual Operator Learning
-To train the parameterized surrogate model $\mathcal{G}_{\theta}^{\Delta}$ of $\mathcal{G}^{\Delta}$ using the residual training set $\mathcal{T}^{res}$, we formulate the following optimization problem to determine the optimal parameter by minimizing the cost function:
-$\ell=\sum_{i=1}^{N}\mathcal{L}(\mathcal{G}_{\theta}^{\Delta}(a_{i},a_{k_{i}}),u_{i}-u_{k_{i}}),$ (10)
-where $\mathcal{L}$ is the cost function for assessing the discrepancy between function residuals.
-The optimization of the physical trajectory residual depicted in Equation 10 can be effectively implemented using existing learning frameworks [2, 11].
-This involves introducing a trajectory residual connection on the explicit output of the base operator network.
-Specifically, we add the auxiliary solution $u_{k_{i}}$ to the predicted residual from the final output of the operator network:
+To mitigate such difficulty, we introduce additional information related to $a_{k_{i}}$ as input:
+  • Taking the corresponding $u_{k_{i}}(x)$ of $a_{k_{i}}(x)$ as the input.
+  • Utilizing calculated relation metrics, e.g. similarity between $a_{i}$ and $a_{k_{i}}$, as inputs.
+Our experimental results (Section˜A.3) indicate that such customized inputs could improve the performance of residual neural operators.
+Physical Residual Connection.
+Based on Equation˜4, to obtain the final predicted solution $\hat{u}_{i}$, we introduce a physical residual connection at the final layer of the base operator network.
+This connection adds the auxiliary solution $u_{k_{i}}$ to the predicted residual:
 $\hat{u}_{i}=Q(v_{l})+u_{k_{i}},$ (13)
-where $Q$ is the last fully connected projector of the operator network, $v_{l}$ is the last latent function in hidden space produced by $\sigma(W_{l}+\mathcal{K}_{l})$ shown in Equation 3.
-During inference, given input function $a^{test}$, we select the auxiliary trajectory $(a_{k_{t}}(x),u_{k_{t}}(x))$ from the training set $\mathcal{T}$ using the sampling function $\mathcal{F}^{ret}$.
-3.2.5 Discussion
-Compared to direct operator learning, residual operator learning presents the following advantages.
-Mitigation of distribution bias.
-Residual operator learning allows us to control the training label distribution, thereby mitigating the data distribution bias.
-We posit that the similarity between input functions and the corresponding output functions’ similarity are positively correlated.
-This correlation is validated empirically in non-chaotic systems, as demonstrated in Darcy Flow experiments (Figure 4).
-Consequently, the distribution of label residuals can be effectively controlled by adjusting the selection range ($K$ in Equation 6) of the sampling function $\mathcal{F}^{ret}$.
-This adjustment allows the construction of sufficiently diverse training labels, thereby alleviating distribution bias in restricted data scenarios and enhancing the generalization capabilities of neural operators.
-Prevent overfitting.
-In scenarios with limited training data, neural networks tend to memorize the training set, leading to significant overfitting.
-Residual learning effectively addresses this issue by modifying the training labels through the use of various auxiliary samples in each training epoch, thus reducing the risk of overfitting.
-Incorporation of explicit physical priors.
-The auxiliary sample acts as a non-parametric physical prior.
-As shown in previous work [12], involving external memory components benefits the model’s generalization capability.
-Moreover, the explicit external memory allows specific needs during inference, such as handling boundary conditions and high resolution samples.
+where $Q$ is the last fully connected projector of the operator network, $v_{l}$ is the last latent function in hidden space produced by $\sigma(W_{l}+\mathcal{K}_{l})$ shown in Equation˜10.
+This enables end-to-end training with existing frameworks while maintaining consistent setups (loss functions, trainable parameters, etc.
+) with the base operator networks.
+Cross-Resolution Sample Integration.
+For cross-resolution scenarios where we need to predict high-resolution outputs $u^{H}$ using low-resolution auxiliary trajectories $(a_{k_{i}},u_{k_{i}})$, we employ Fourier-based upsampling to align the resolutions before integration:
+$a_{input}=\text{UpSampling}(a_{k_{i}})\oplus a^{H}.$ (14)
+Similarly, for the residual connection, we upsample the low-resolution auxiliary solution:
+$\hat{u}^{H}=Q(v_{l})+\text{UpSampling}(u_{k_{i}}),$ (15)
+where $Q$ is the decoder and $v_{l}$ is the final hidden state.
+The upsampling is performed through zero-padding in Fourier space to preserve the low-frequency structures while extending to higher resolution, formulated as follows:
+$\text{UpSampling}(a)=\mathcal{F}_{\text{fourier}}^{-1}(\text{ZeroPad}(\mathcal{F}_{\text{fourier}}(a))),$ (16)
+where $\mathcal{F}_{\text{fourier}}$ and $\mathcal{F}_{\text{fourier}}^{-1}$ denote the Fourier transform and its inverse, and ZeroPad extends the spectral representation with zeros to match the target resolution.
+This Fourier-based approach ensures smooth upsampling while preserving the physical characteristics of the auxiliary trajectories.
+3.5 How DeltaPhi Improves Generalization Capability
+In data-limited scenarios, DeltaPhi enhances generalization by mitigating overfitting, which is achieved through a form of implicit data augmentation  [^1] We name this implicit data augmentation because there are no more explicit input-output physical function pairs, but directly training on more diverse residual pairs.
+enabled by our residual learning framework.
+Mitigating Overfitting via Implicit Data Augmentation.
+The core benefit of residual learning is its effectiveness in mitigating overfitting, particularly in data-limited settings.
+This is achieved through two complementary mechanisms.
+First, DeltaPhi functions as an implicit data augmentation strategy.
+For a training set with $N$ samples, direct learning provides only $N$ input-output pairs.
+In contrast, DeltaPhi transforms each sample $(a_{i},u_{i})$ into $K$ distinct residual pairs $\{(a_{i},a_{k}),u_{i}-u_{k}\}^{K}_{k=1}$ by pairing it with $K$ different auxiliary samples selected via Section˜3.3.
+The effectiveness of this augmentation stems from the stability property of PDEs (Section˜3.1), which establishes a structured relationship between input differences and solution residuals.
+This property allows us to meaningfully control the diversity of the augmented training residuals by sampling auxiliary inputs with varying similarity levels.
+Therefore, with an opportune retrieval range, we can expand the sparse training distribution—thus mitigating distribution bias—while ensuring the residual learning task remains feasible and does not involve highly dissimilar, difficult-to-learn pairs.
+This process significantly increases the density and diversity of the training distribution, directly addressing the problem of sparse training data.
+Second, the residual learning formulation itself discourages memorization.
+Because the predicted residual $\mathcal{G}_{\theta}^{\Delta}(a_{i},a_{k_{i}})$ must adapt to different auxiliary samples $a_{k_{i}}$ for the same input $a_{i}$, the model is discouraged from simply memorizing a fixed input-output mapping.
+Thus, the model is less prone to memorizing labels for specific input functions in data-limited scenarios.
+From a geometric perspective, this framework effectively shifts the learning paradigm from approximating isolated points (absolute solutions) on the solution manifold to learning tangent vectors (solution residuals) that describe the manifold’s local geometry.
+Learning these relational vectors provides a richer learning signal.
+The stability property discussed in Section˜3.1 ensures this task is tractable, as it establishes a structured relationship between input differences and solution residuals, allowing us to expand the training distribution while maintaining learnability.
+Tractability of Residual Mapping Learning.
+It’s noted that learning the residual mapping $\mathcal{G}^{\Delta}$ can be more difficult than learning the direct mapping $\mathcal{G}$.
+However, this increased difficulty does not arise from the introduction of spurious high-frequency noise; if two solution functions are smooth, their residual will also be smooth.
+Instead, the challenge stems from the inherent complexity of the residual mapping itself, which requires the model to synthesize information from two separate input functions ($a_{i}$ and $a_{k_{i}}$).
+Despite this increased complexity, our experiments demonstrate that existing neural operators are fully capable of learning this mapping effectively, even learning to respect its inherent symmetry without explicit constraints (see Section˜A.8).
+Influence of Retrieval Range.
+While the retrieval range controls the diversity of residuals for training, our results in Section˜A.1 show DeltaPhi is robust to the selection of this hyperparameter.
+For experimental consistency, we set $K=20$ throughout our comparisons.
+We further analysis its impact across different test splits in Section˜A.2 and discuss its selection in Section˜C.1.
+In addition, we introduce an alternative sampling strategy without $K$ selection in Section˜A.6.
 4 Experiment
-This section provides the experiments on various PDEs, resolution generalization problem, and statistical analysis of residual operator learning.  Appendix A presents additional ablation studies.
-4.1 Residual Learning vs. Direct Learning on Various PDEs
-This section validates residual operator learning on various PDEs.
-The experiment covers both irregular domain problems and regular domain problems.
-In addition, we test the performance improvement under variant training data amount.
-To ensure fair comparisons, all baselines and our counterpart models are compared with exactly same hyperparameters, trainable weight amount, optimizer configuration, and training step.
+Figure 2: Prediction error visualization on irregular domain problems.
+This section provides the comparison on various PDEs, comparison on resolution generalization problems, and statistical analysis of residual operator learning.
+We present more experimental results and analysis in Appendix˜A.
+4.1 Comparison on Various PDEs
+We evaluate residual operator learning across irregular and regular domain PDEs, testing performance with different training data sizes.
+For fair comparisons, all comparisons use exactly identical hyperparameters, model weights, optimizer settings, and training steps.
 Irregular domain problems.
-We validate the effectiveness of the proposed approach on irregular domain problems.
-The experimental setup is exactly same as [10].
-We experiment with five problems covering both 2D and 3D scenarios, i.e.
-Irregular Darcy Flow (2D), Pipe Turbulence (2D), Heat Transfer (3D), Composite (3D), and Blood Flow (3D), all of which are defined on irregular physical domains and processed as a set of discrete triangle meshes.
-The baselines contain GraphSAGE [43], DeepOnet [3], POD-DeepOnet [44], FNO [2] and NORM [10].
-We implement the residual learning version of NORM and compare it with the direct learning model under exactly same setup.
-Appendix B presents more experimental details about the datasets and baselines.
-We report the results in Table 1.
-The residual neural operator NORM (DeltaPhi) consistently outperforms all baselines on all irregular domain problems.
-The enhancement is particularly significant on Pipe Turbulence and Heat Transfer problem, with 40% - 50% relative improvement (the row "Relative Gain" in Table 1) over original NORM, i.e.
-the NORM (Previous).
-These results indicate that the proposed residual operator learning works well on irregular domain problems.
-Figure 2: Prediction error visualization on irregular domain problems.
-In addition, we visualize the prediction error in Figure 2.
-The error value is calculated as the absolute difference between the predicted function $\hat{u}$ and ground-truth function $u^{GT}$, i.e.
-$|\hat{u}-u^{GT}|$.
-Compared to previous direct learning methods, DeltaPhi significantly reduces the error scale.
+Following (chen2023learning, ), we evaluate five problems: 2D (Irregular Darcy Flow, Pipe Turbulence) and 3D (Heat Transfer, Composite, Blood Flow) scenarios, all defined on irregular domains represented by discrete triangle meshes.
+We compare against GraphSAGE (hamilton2017inductive, ), DeepOnet (lu2019deeponet, ), POD-DeepOnet (lu2022comprehensive, ), FNO (li2020fourier, ) and NORM (chen2023learning, ) baselines, implementing a residual learning version of NORM for direct comparison.
+Appendix˜B provides more details.
+Results in Table˜1 show that NORM (DeltaPhi) outperforms all baselines across irregular domain problems.
+Performance gains are most notable for Pipe Turbulence and Heat Transfer, showing 40-50% improvement over NORM, demonstrating the residual operator’s effectiveness.
 Table 1: Relative error comparison of operator learning on irregular domains.
@@
 FNO 3.83e-2 3.80e-2 - - -
-NORM (Previous) 1.05e-2 1.01e-2 2.70e-4 9.99e-3 4.82e-2
+NORM (Direct) 1.05e-2 1.01e-2 2.70e-4 9.99e-3 4.82e-2
 NORM (DeltaPhi) 9.75e-3 5.96e-3 1.35e-4 9.18e-3 4.29e-2
 Relative Gain 7.14% $\uparrow$ 40.99% $\uparrow$ 50.00% $\uparrow$ 8.11% $\uparrow$ 11.00% $\uparrow$
+In addition, we visualize the prediction error in Figure˜2.
+The error value is calculated as the absolute difference between the predicted function $\hat{u}$ and ground-truth function $u^{GT}$, i.e.
+$|\hat{u}-u^{GT}|$.
+Compared to previous direct learning methods, DeltaPhi significantly reduces the error scale.
 Regular domain problems.
 We conduct the experiment on two regular domain problems i.e.
-Darcy Flow and Navier-Stokes, following [2].
+Darcy Flow and Navier-Stokes, following (li2020fourier, ).
 The training data amount is 100.
-We validate DeltaPhi on different neural operators, including FNO [2], FFNO [11], CFNO [16], GNOT [23], Galerkin [34], MiOnet [45] and Resnet [42].
-The performance is measured with relative error.
-Appendix B presents more experimental details.
-Table 2 presents the experimental results.
-On both Darcy Flow and Navier-Stokes Equations, the proposed residual operator learning (DeltaPhi) consistently improves performance across different base models.
-The results validate the effectiveness of DeltaPhi on regular domain problems.
-However, we observe that the performance improvement of the model on the Navier-Stokes Equation is relatively marginal.
-We conjecture this is due to the complexity of the fluid dynamic with the viscosity $\upsilon=1e-5$, whereby the correlation between input and output functions with long time intervals is diminished, i.e.
-similar initial functions $a(x)$ do not necessarily correspond to sufficiently similar output functions $u(x)$.
-This is a limitation of the current implementation that does not affect the conclusion of this work.
-It could be alleviated with additional designs such as performing auxiliary trajectory retrieval at each step of time-series prediction to mitigate the decay in correlation between input and output functions.
+We validate DeltaPhi on different neural operators, including FNO (li2020fourier, ), FFNO (tran2021factorized, ), CFNO (brandstetter2022clifford, ), GNOT (hao2023gnot, ), Galerkin (cao2021choose, ), MiOnet (jin2022mionet, ) and Resnet (he2016deep, ).
+The metric is relative L2 error.
+Appendix˜B presents more details.
+Table˜2 shows DeltaPhi improves performance across all base models for both Darcy Flow and Navier-Stokes equations, confirming its effectiveness on regular domains.
+The improvement on the Navier-Stokes, while modest compared to Darcy Flow, is particularly noteworthy given the inherently chaotic nature of fluid dynamics at viscosity $\upsilon=1e-5$.
+Even within this challenging scenario characterized by weakened correlation between similar initial conditions $a(x)$ and their evolved states $u(x)$, DeltaPhi still achieves consistent gains.
+This robustness to complex dynamics strengthens conclusions regarding the method’s general applicability.
+Performance on temporal systems could be further enhanced through designs such as performing auxiliary sample retrieval at each step of time-series prediction to mitigate the decay in correlation between input and output functions.
 Table 2: Relative error comparison on regular Darcy Flow and Navier-Stokes Equation.
-Model Darcy Flow Navier-Stokes Equation
+Model
+Darcy Flow Navier-Stokes Equation
 Direct Learning
@@
 Resnet 1.25e-1 1.07e-1 14.14% $\uparrow$ 4.39e-1 4.30e-1 1.87% $\uparrow$
-Table 3: Relative error comparison on different training scales.
-Model Different Train Size
+Table 3: Comparison on different training scales.
+Model
+Different Train Size
 100
@@
 900
-FNO (Previous) 3.70e-2 1.40e-2 1.02e-2 8.37e-3 7.39e-3
+FNO (Direct) 3.70e-2 1.40e-2 1.02e-2 8.37e-3 7.39e-3
 FNO (DeltaPhi) 3.31e-2 1.34e-2 9.64e-3 8.06e-3 7.18e-3
 Relative Gain 10.54% $\uparrow$ 4.08% $\uparrow$ 5.49% $\uparrow$ 3.74% $\uparrow$ 2.83% $\uparrow$
-FFNO (Previous) 5.22e-2 1.69e-2 9.73e-3 7.48e-3 6.16e-3
+FFNO (Direct) 5.22e-2 1.69e-2 9.73e-3 7.48e-3 6.16e-3
 FFNO (DeltaPhi) 2.93e-2 1.11e-2 7.20e-3 6.22e-3 5.30e-3
 Relative Gain 43.76% $\uparrow$ 34.47% $\uparrow$ 25.98% $\uparrow$ 16.79% $\uparrow$ 14.02% $\uparrow$
-Resnet (Previous) 1.25e-1 1.22e-1 1.21e-1 1.21e-1 1.21e-1
-Resnet (DeltaPhi) 1.07e-1 9.42e-2 9.08e-2 8.63e-2 8.31e-2
-Relative Gain 14.14% $\uparrow$ 22.74% $\uparrow$ 25.14% $\uparrow$ 28.92% $\uparrow$ 31.51% $\uparrow$
 Different training scales.
-We test the residual operator learning on different training set sizes.
-The experiment is conducted on Darcy Flow problems, with three base models FNO, FFNO, and Resnet.
-Specifically, we train all models with five training data scales 100, 300, 500, 700, and 900 respectively, and evaluate their performance on the additional same 100 testing samples.
-The experimental results are shown in Table 3.
+We train residual operators using FNO, FFNO, and Resnet on Darcy Flow with varying training set sizes (100, 300, 500, 700, and 900) and test their performance on the additional 100 samples.
+The experimental results are shown in Table˜3.
 On different training set sizes, DeltaPhi improves the prediction performances of neural operators.
@@
 In addition, several residual neural operators even outperform direct neural operators using less training data amount, e.g.
-FFNO (DeltaPhi) with training size 500 outperforms FFNO (Previous) with training size 700, Resnet (DeltaPhi) with training size 700 outperforms Resnet (Previous) with training size 900.
-The experimental results empirically demonstrate the effectiveness of DeltaPhi across different training set scales.
-4.2 Residual Learning vs. Direct Learning on Resolution Generalization Problem
+FFNO (DeltaPhi) with training size 500 outperforms FFNO (Direct) with training size 700, Resnet (DeltaPhi) with training size 700 outperforms Resnet (Direct) with training size 900.
+The results empirically demonstrate the effectiveness of DeltaPhi across different training set scales.
+4.2 Comparison on Resolution Generalization Problem
 This section presents the experimental results of training-free resolution generalization.
 Resolution generalization problem.
-Fourier neural operator [2] enables zero-shot resolution generalization inference.
+Fourier neural operator (li2020fourier, ) enables zero-shot resolution generalization inference.
 However, when the training data grid is excessively coarse, the inference performance over high resolution data drops a lot.
@@
 Setup.
-We conduct the training-free resolution generalization experiment on Darcy Flow [2].
-Specifically, we train the models with low resolutions $85\times 85$, $43\times 43$, $31\times 31$, and $22\times 2$ 22 respectively, then evaluate them using high resolution $421\times 421$ test data.
-Both FNO [2] and FFNO [11] are compared as the base models.
+We conduct the training-free resolution generalization experiment on Darcy Flow (li2020fourier, ).
+Specifically, we train the neural operators with low resolutions $85\times 85$, $43\times 43$, $31\times 31$, and $22\times 22$ respectively, then evaluate them using high resolution $421\times 421$ test data.
+Both FNO (li2020fourier, ) and FFNO (tran2021factorized, ) are compared as the base models.
 The training set scale includes 100 and 900 trajectories.
 Table 4: Relative error comparison on zero-shot resolution generalization.
-Model Train Size=100 Train Size=900
+Model
+Train Size=100 Train Size=900
 $\mathbb{85\times 85}$
@@
 $\mathbb{22\times 22}$
-FNO (Previous) 7.29e-2 1.19e-1 1.49e-1 1.70e-1 6.73e-2 1.11e-1 1.34e-1 1.76e-1
+FNO (Direct) 7.29e-2 1.19e-1 1.49e-1 1.70e-1 6.73e-2 1.11e-1 1.34e-1 1.76e-1
 FNO (DeltaPhi) 6.91e-2 1.13e-1 1.33e-1 1.53e-1 4.91e-2 9.37e-2 1.15e-1 1.57e-1
 Relative Gain 5.13% $\uparrow$ 5.06% $\uparrow$ 10.43% $\uparrow$ 10.08% $\uparrow$ 27.04% $\uparrow$ 15.65% $\uparrow$ 14.29% $\uparrow$ 10.31% $\uparrow$
-FFNO (Previous) 6.64e-2 1.04e-1 1.24e-1 1.43e-1 4.89e-2 1.01e-1 1.27e-1 1.49e-1
+FFNO (Direct) 6.64e-2 1.04e-1 1.24e-1 1.43e-1 4.89e-2 1.01e-1 1.27e-1 1.49e-1
 FFNO (DeltaPhi) 6.34e-2 9.90e-2 1.18e-1 1.36e-1 4.42e-2 9.00e-2 1.12e-1 1.35e-1
@@
 Performance comparison.
-Table 4 presents the resolution generalization results.
-Compared to FNO [2], FFNO [11] shows prefer generalization ability for its less number of network weights.
+Table˜4 presents the resolution generalization results.
+Compared to FNO (li2020fourier, ), FFNO (tran2021factorized, ) shows preferable generalization ability for its less number of network weights.
 In all experimental settings, the proposed physical residual models (tagged with "DeltaPhi") evidently outperform their counterpart base models.
 The performance improvement of residual operator learning is more noticeable when the training set size is 900.
-In this setting, FNO (DeltaPhi) managed to match, or even surpass FFNO [11] on resolution $85\times 85$, $43\times 43$, and $22\times 2$ 22, despite using the relatively weaker base network.
+In this setting, FNO (DeltaPhi) managed to match, or even surpass FFNO (tran2021factorized, ) on resolution $85\times 85$, $43\times 43$, and $22\times 22$, despite using the relatively weaker base network.
 The results empirically draw the conclusion that physical residual learning improves zero-shot resolution generalization performance.
 Training curve comparison.
-We also report the test loss curve during training in  Figure 3.
-The training data is 900 and the training resolution is $85\times 85$.
+We also report the test loss curve during training in  Figure˜3.
+The training amount is 900 and the training resolution is $85\times 85$.
 As the training proceeds, the relative error on resolution $85\times 85$ data continues to decline.
 However, after some epochs, the loss on resolution $421\times 421$ gradually increases due to overfitting.
-The proposed physical residual operator (named with DP-FNO) significantly mitigates this overfitting tendency of the base model (named with FNO).
+The proposed residual operator (named with DP-FNO) significantly mitigates this overfitting tendency of the base model (named with FNO).
 4.3 Statistical Analysis of Residual Operator Learning
-This section further validates the generalization property of DeltaPhi through statistical analysis.
+This section validates DeltaPhi’s foundations through statistical analysis.
+Figure 4: The curve of distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as retrieval similarity rank increases.
 Similarity between output functions.
-We statistically validate the assumption (mentioned in Section 3.2.5) that the similarities between the input function pair and the corresponding output function pair are positively related.
-This means that for input function $a^{test}(x)$, the higher similarity score of the auxiliary sample $(a_{k_{t}}(x),u_{k_{t}}(x))$, the lower difference between the targeted solution function $u^{test}(x)$ and auxiliary solution function $u_{k_{t}}(x)$.
-Figure 4: The variation curve of the normalized distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as retrieval similarity rank increases. Different color corresponds to different retrieval set sizes.
-In Figure 4, we visualize the variation curve of the normalized distance between $u^{test}(x)$ and $u_{k_{t}}(x)$ as the similarity rank of $a_{k_{t}}(x)$ increases on Darcy Flow.
-(Section D.1 presents the visualization details.) The normalized distance consistently increases as the similarity rank increases across different retrieval set size.
-This result concludes that the similarity between output function pair is positively correlated with the similarity between input function pair, thus the training label distribution could be effectively controlled for alleviating the training label distribution bias.
+We validate the stability property in Section˜3.1, which states that similar inputs lead to similar solutions.
+This property is essential for our residual learning approach.
+In Figure˜4, we show how the normalized distance between output functions $u^{test}(x)$ and $u_{k_{t}}(x)$ changes as the similarity rank of their input functions increases for Darcy Flow.
+(Section˜D.1 provides visualization details.) The normalized distance grows consistently with similarity rank across all retrieval set sizes.
+This confirms the stability property - when input functions become less similar, their solutions also become less similar.
+The relationship also shows that increasing the retrieval range $K$ during training adds more diversity to the training samples by including a wider range of residual patterns, thereby expanding the learned solution space while maintaining physical validity.
+Figure 5: Label distribution visualization.
+The points represent dimension-reduced labels ($u_{*}(x)$ for direct learning, $u_{*}(x)-u_{k_{*}}(x)$ for residual learning) through Principal Component Analysis.
+The ellipses with dotted lines and solid lines represent standard deviation and range, respectively.
+Green and red color correspond to training and testing set, respectively.
 Training label distribution comparison.
-Through training label distribution visualization, we statistically validate the enhanced generalization property of DeltaPhi on Darcy Flow.
-We process both the training dataset and the testing dataset with the following steps.
-First, the high-dimensional labels (output function) are compressed into 2 dimensions.
-Then we plot the 2D point of every label and the circles of standard deviation and range for all samples.
-Section D.2 presents more details.
-Figure 5: Label distribution visualization on Darcy Flow.
-The points represent dimension-reduced labels ($u_{*}(x)$ for direct learning, $u_{*}(x)-u_{k_{*}}(x)$ for residual learning) through Principle Component Analysis.
-The ellipses with dotted lines and solid lines represent the standard deviation and range, respectively.
-Green and red color correspond to training set and testing set, respectively.
-Figure 5 presents the visualization results.
-(1) As depicted in the left part of Figure 5, in direct operator learning the restricted amount of training data leads to the limited diversity of training labels.
-Consequently, it is challenging to comprehensively cover the distribution of testing labels with training labels.
-We refer to this situation as training label distribution bias.
-In such scenarios, the trained neural operator models struggles with managing certain solutions to be solved.
-(2) As depicted in the right part of Figure 5, compared to direct operator learning, the testing label distribution becomes more concentrated, and the training labels distribution becomes more dispersed in residual operator learning.
-The distribution bias between training labels and testing labels is mitigated as mentioned in Section 3.2.5.
+We demonstrate how DeltaPhi provides implicit data augmentation on Darcy Flow by visualizing label distributions.
+We project both training and testing labels (output functions for direct learning, output residuals for DeltaPhi) into 2D using PCA.
+Section˜D.2 details this process.
+Figure˜5 reveals the main benefits of our residual learning framework: In direct operator learning (left), limited training data restricts coverage of the solution space, making generalization difficult.
+This shows the fundamental challenge of distribution bias in data-limited scenarios.
+With DeltaPhi (right), we see that: (a) the testing label distribution becomes more concentrated, showing that learning residuals is easier than learning absolute solutions, and (b) the training labels show greater spread, demonstrating the implicit data augmentation claimed in Section˜3.5.
+By creating diverse valid residual pairs through controlled auxiliary sample retrieval, DeltaPhi effectively enhances the training distribution while preserving physical validity.
+5 Limitation
+Despite the comprehensive validation of DeltaPhi’s superiority through an extensive range of experiments, we recognize some inherent limitations that, nevertheless, do not impact the solidity of our conclusions.
+First, although highly challenging, it is significant to study more realistic applications where training data acquisition is extremely costly, such as high-fidelity fluid modeling or biomedical simulations.
+Second, we acknowledge the potential constraints of using a general cosine similarity metric.
+While effective, this metric may not be optimal for all physical systems.
+This is particularly relevant for highly chaotic systems, such as the low-viscosity Navier-Stokes equations (Table˜2), where DeltaPhi, while effective, showed a more modest relative improvement compared to other problems.
+This suggests that the complexities associated with such systems, where simple initial state similarity may not sufficiently capture the relationship between solution trajectories, warrant further investigation.
+Future work could explore these aspects, for instance, by developing more advanced auxiliary sample retrieval strategies better suited for complex dynamics, which may unlock greater performance gains in these challenging applications.
 6 Conclusion
-This work proposes to learn physical trajectory residuals for PDE solving based on existing neural operator networks.
-Extensive experiments substantiate that learning residual operator mapping is preferred compared to previous learning direct operator mapping between input-output function fields.
-We hope the introduced method provides insights for future development of machine learning based PDE solving or other high-dimensional regression problems.
-In future work, we will explore employing DeltaPhi in practical scenarios of data deficiency and bias, e.g.
-realistic physical phenomenon forecasting, or high-fidelity physics dynamic modeling.
-6 Limitation
-Despite the comprehensive validation of the superiority of Physical Residual Operator Learning through an extensive range of experiments, we recognize some inherent limitations that, nevertheless, do not impact the solidity of our conclusions.
-First, when the correlation between the input and output functions of PDEs is weak, such as in chaotic systems, the performance enhancement of the current implementation may be constrained.
-Second, compared to direct operator learning, the proposed residual operator learning increases the learning difficulty, thus requiring the base model to have sufficient representation capability.
-Although beyond the scope of this work, these aspects warrant further exploration in future studies.
-Appendix A Ablation Studies
-A.1 Influence of Similar Trajectory Retrieval
-In this section, we test the impact of different retrieval ranges $K$ presented in Equation 6 under various training data scales.
+This work proposes to learn physical state residuals for PDE solving by reformulating the task from direct mapping to learning residuals between similar physical states, supported by physical system stability.
+This approach enables implicit data augmentation without requiring additional data collection, effectively addressing challenges in data-limited scenarios.
+We validate the effectiveness across various physical systems, neural architectures, and both regular and irregular domains.
+We hope this framework provides insights for future development of machine learning based PDE solving.
+Acknowledgments
+This work is partially supported by National Science and Technology Major Project (2022ZD0117802).
+This work was supported in part by General Program of National Natural Science Foundation of China (62372403) and "Pioneer" and "Leading Goose" R&D Program of Zhejiang (No.
+2025C02032).
+This work is also supported by the Fundamental Research Funds for the Central Universities (226-2025-00080) and the Earth System Big Data Platform of the School of Earth Sciences, Zhejiang University.
+Appendix A Additional Experiment
+A.1 Impact of Retrieval Range and Random Sampling
+This section analyzes how different retrieval range values ($K$) affect the performance of our residual learning framework.
+Additionally, we investigate the importance of similarity-based retrieval by replacing it with random sampling (denoted as "w/o $\mathcal{F}^{\text{ret}}$").
+Table 5: Relative error comparison of FNO (DeltaPhi) with different retrieval ranges ($K$) and training set sizes on Darcy Flow.
+The symbols $\uparrow$ and $\downarrow$ denote performance increasing and decreasing respectively, compared to direct learning.
+Training
+Set Size
+FNO
+FNO (DeltaPhi) with different $K$ values w/o $\mathcal{F}^{\text{ret}}$
+KK=5 KK=10 KK=20 KK=30 KK=40 KK=50
+100 3.57e-2 3.34e-2 $\uparrow$ 3.31e-2 $\uparrow$ 3.31e-2 $\uparrow$ 3.31e-2 $\uparrow$ 3.36e-2 $\uparrow$ 3.30e-2 $\uparrow$ 3.04e-2 $\uparrow$
+300 1.41e-2 1.38e-2 $\uparrow$ 1.41e-2 $\uparrow$ 1.35e-2 $\uparrow$ 1.36e-2 $\uparrow$ 1.40e-2 $\uparrow$ 1.37e-2 $\uparrow$ 1.62e-2 $\downarrow$
+500 1.02e-2 1.01e-2 $\uparrow$ 1.01e-2 $\uparrow$ 9.64e-3 $\uparrow$ 9.96e-3 $\uparrow$ 1.00e-2 $\uparrow$ 1.01e-2 $\uparrow$ 1.29e-2 $\downarrow$
+700 8.37e-3 8.29e-3 $\uparrow$ 8.55e-3 $\downarrow$ 8.06e-3 $\uparrow$ 8.04e-3 $\uparrow$ 8.03e-3 $\uparrow$ 8.19e-3 $\uparrow$ 1.10e-2 $\downarrow$
+Impact of retrieval range $K$. Table˜5 presents the relative error of FNO (DeltaPhi) with different retrieval range values across various training set sizes. We observe several important patterns:
+  • In data-limited scenarios (100 to 500 samples), DeltaPhi demonstrates robust performance across different $K$ values, obtaining consistent performance gains across diverse selection of $K$.
+  • With very limited training data (100 samples), larger $K$ values (40-50) tend to perform slightly better, suggesting that expanding the retrieval range becomes beneficial when the training data is extremely sparse.
+  • Conversely, with abundant training data (700 samples), moderate $K$ values (20-40) generally outperform both smaller and larger retrieval ranges, indicating an optimal balance between sample diversity and learning difficulty.
+The consistent performance across different $K$ values demonstrates that our framework is not overly sensitive to this hyperparameter, making it practical for real-world applications where exhaustive hyperparameter tuning may not be feasible.
+Importance of similarity-based retrieval.
+The last column of Table˜5 shows the performance when auxiliary samples are randomly selected from the training set without considering similarity ("Random Sampling").
+Comparing this with similarity-based retrieval reveals:
+  • For most training set sizes (700, 500, and 300), random sampling performs significantly worse than similarity-based retrieval, with error increases compared to the best $K$ value configuration.
+  • This performance degradation confirms our discussion in Section˜3.1: the stability property of PDEs makes learning residuals between similar states more tractable than between arbitrary states.
+These results empirically validate the importance of similarity-based auxiliary sample retrieval in our framework, particularly as the training set size increases.
+The optimal retrieval strategy may depend on the specific characteristics of the PDE system and the available training data, but similarity-based retrieval provides a robust default approach for most scenarios.
+A.2 Performance Analysis on Different Test Distribution
+To evaluate the robustness of DeltaPhi under varying test distributions, we conduct experiments on test samples with different distances from the training set.
+Specifically, we first compute the cosine similarity score between each test sample and its closest training sample on Darcy Flow.
+Based on these scores, we divide the test set into 7 non-overlapping splits, where a lower average score indicates a more challenging split due to greater deviation from the training distribution.
+Table 6: Statistics of different test splits based on similarity to training data.
+Sample Number Min Score Max Score Average Score
+Split1 23 0.85 0.86 0.86
+Split2 14 0.86 0.88 0.87
+Split3 30 0.88 0.90 0.89
+Split4 37 0.90 0.91 0.90
+Split5 33 0.91 0.93 0.92
+Split6 43 0.93 0.95 0.94
+Split7 16 0.95 0.96 0.95
+We evaluate both direct operator learning and residual operator learning using FNO as the base model across these splits.
+Additionally, we test different retrieval range settings ($K$) to analyze their impact on performance.
+The results are presented in Table˜7.
+Table 7: Relative error comparison on different test splits. Numbers in parentheses show relative improvement over direct learning.
+Methods Split1 Split2 Split3 Split4 Split5 Split6 Split7
+Direct Learning 5.08e-2 5.00e-2 4.71e-2 3.31e-2 2.77e-2 2.73e-2 2.43e-2
+DeltaPhi (KK=20) 4.77e-2 4.68e-2 4.32e-2 3.14e-2 2.58e-2 2.46e-2 2.18e-2
+[…]
+DeltaPhi (KK=50) 4.80e-2 4.66e-2 4.32e-2 3.13e-2 2.57e-2 2.47e-2 2.11e-2
+[…]
+DeltaPhi (KK=5) 4.83e-2 5.01e-2 4.39e-2 3.27e-2 2.52e-2 2.43e-2 2.10e-2
+[…]
+The experimental results reveal several key findings:
+  • DeltaPhi with appropriate retrieval ranges (KK=20 or KK=50) consistently outperforms direct operator learning across all splits, even on the most challenging ones (Split1, Split2, Split3) that significantly deviate from the training distribution.
+  • The choice of retrieval range $K$ has a notable impact on performance.
+A relatively small $K$ value (KK=5) leads to smaller improvements on challenging splits (Split1, Split2, and Split4), and can even underperform direct learning in some cases (Split2).
+  • $K$ values (KK=20 or KK=50) in appropriate ranges generally provide stable improvements across different splits, suggesting they enable more robust generalization under distribution shifts.
+These results demonstrate that DeltaPhi maintains its effectiveness even when test samples significantly differ from training data, supporting its robustness in practical scenarios where training and testing distributions may not perfectly align.
+A.3 Customized Auxiliary Input Ablation
+This section conducts ablation experiments on the customized auxiliary input.
 Setup.
-The experimental PDEs are Darcy Flow [2] and Navier-Stokes [2].
-We randomly sample 100, 300, 500, 700, and 900 data for training, and evaluate the trained models on an additional 200 test data with the same resolution.
-FNO [2] and Resnet [42] are included as the base model for comparison.
-Besides, we experiment with varying retrieval range $K$ on different training data scales and evaluate the performance without utilizing Similar Trajectory Retrieval.
-(a) Darcy FLow
-(b) Navier-Stokes
-Figure 6: Influence of similar trajectory retrieval range $K$ (presented in Equation 6). "DP-*" denotes the proposed residual learning version of each base model.
-Performance Comparison.
-The experimental results are shown in  Figure 6.
-In most settings, the physical trajectory residual learning framework DP-FNO and DP-Resnet outperform their corresponding baselines FNO [2] and Resnet [42] respectively.
-In the small data regime, this performance enhancement is particularly evident.
-Additionally, on Darcy Flow in Figure 6(a), as the amount of training data increases, the performance increasing of the weak base model Resnet [42] gradually converges, while the corresponding residual operator DP-Resnet continually improves.
-Same as Table 3, the experimental results validate that the proposed residual operator learning performs better across various training set scales, especially on small data scenarios.
-Influence of Similar Trajectory Retrieval.
-Figure 6 presents the physical residual learning with varying similar trajectory retrieval range $K$, as well as the simply random sampling without considering similarity (could also be seen as setting $K$ as the training set size).
-On various settings, as the sampling similarity range $K$ increases, the overall trend is that the model’s performance first rises and then falls.
-This suggests that setting an appropriate range $K$ is of significant importance.
-Another point worth noting is that, in nearly all settings, the model’s performance appears to be the poorest when similarity-based retrieval is removed.
-This demonstrates the crucial role of Similar Trajectory Retrieval.
-These experimental results lead to the conclusion that Similar Trajectory Retrieval plays an important role in the physical residual learning framework.
-A.2 Customized Auxiliary Input Ablation
-This section conducts ablation experiments on the customized auxiliary input.
+We evaluate the influence of customized auxiliary input on Navier-Stokes.
+The training trajectory number is 100.
+FNO [2] is used as the base model.
+We test 3 types input of $a_{k_{i}}(x)$, including complete $a_{k_{i}}(x)$, empty $a_{k_{i}}(x)$ and partial (only input last 3 trajectory steps) $a_{k_{i}}(x)$.
+In addition, we experiment with the inclusion or omission of $u_{k_{i}}(x)$, as well as the addition or exclusion of $Score_{i,k_{i}}$.
 Table 8: Ablation of customized auxiliary input on Navier-Stokes. Bold font, underline and wavyline respectively denote the best, second best, and inferior results.
-Customized Auxiliary Input Relative Error
+Customized Auxiliary Input
+Relative Error
 $a_{k_{i}}(x)$ $u_{k_{i}}(x)$ $Score_{i,{k_{i}}}$
@@
 Partial 2.37e-1
-Setup.
-We evaluate the influence of customized auxiliary input on Navier-Stokes.
-The training trajectory number is 100.
-FNO [2] is used as the base model.
-We test 3 types input of $a_{k_{i}}(x)$, including complete $a_{k_{i}}(x)$, empty $a_{k_{i}}(x)$ and partial (last 3 trajectory steps) $a_{k_{i}}(x)$.
-In addition, we experiment with the inclusion or omission of $u_{k_{i}}(x)$, as well as the addition or exclusion of $Score_{i,k_{i}}$.
 Customized Auxiliary Input Influence.
-The results are shown in Table 5.
+The results are shown in Table˜8.
 The first three rows indicate that utilizing partial $a_{k_{i}}(x)$ yields the best results for $a_{k_{i}}(x)$.
@@
 The above findings suggest that the utilization of customized input is beneficial, thereby confirming the conclusion that proper customized input could improve performance.
-A.3 Auxiliary Trajectory Selection during Inference
+A.4 Auxiliary Sample Selection during Inference
+As described in Section˜3.3, we greedily select take the auxiliary sample with the top similarity score during inference.
+To validate the robustness of the trained residual neural operator on different auxiliary samples, we experiment with randomly sampled auxiliary trajectories from the top-10 similarity scores.
+The experiment is conducted on Darcy Flow with FNO as the base model.
 Table 9: Results of repeat Evaluation using random auxiliary sample with Top-10 similarity score.
@@
 STD 5.40649e-5
-As described in Section 3.2.2, we greedily select take the auxiliary sample with the top similarity score during inference.
-To validate the robustness of the trained residual neural operator on different auxiliary samples, we experiment with randomly sampled auxiliary trajectories from the top-10 similarity scores.
-The experiment is conducted on Darcy Flow with FNO as the base model.
-Table 6 presents the relative error of every experimental run and the standard deviation (STD) across five repeated experiments.
+Table˜9 presents the relative error of every experimental run and the standard deviation (STD) across five repeated experiments.
 The low standard deviation value 5.41e-5 indicates that the trained residual neural operators are not sensitive to the selection of auxiliary samples.
 This concludes that the neural operator effectively learns the proposed residual operator mapping.
-A.4 Different Similarity Function
-Table 7: Results of DeltaPhi-FNO using different similarity function.
+A.5 Different Similarity Function
+We use Cosine Similarity for its two advantages:
+  • Unaffected by the magnitude of the numerical value. Thus it could be conveniently applied to various PDEs without considering the numerical scale distribution of physical fields.
+  • Taking normalized values in $[0,1]$.
+We could simply use it as an additional network input.
+As Table˜8 shows, incorporating the similarity scores improves the model’s performance on the Navier-Stokes equation.
+To explore more retrieval metrics for scientific machine learning, we experiment with other discrepancy metrics (all the distance values are calculated after normalization), as shown in Table˜10.
+Table 10: Results of DeltaPhi-FNO using different similarity functions.
 Similarity Functions Relative Error
@@
 Manhattan Distance 3.28e-2
-We use Cosine Similarity for its two advantages:
-  • Unaffected by the magnitude of the numerical value. Thus it could be conveniently applied to various PDEs without considering the numerical scale distribution of physical fields.
-  • Taking normalized values in $[0,1]$.
-We could simply use it as an additional network input.
-As Table 5 shows, incorporating the similarity scores improves the model’s performance on the Navier-Stokes equation.
-To explore more retrieval metrics for scientific machine learning, we experiment with other discrepancy metrics (all the distance values are calculated after normalization), as shown in Table 7.
 The results indicate that different distance functions yield subtle impacts.
@@
 Additionally, inspired by the great success of the NLP community, conducting retrieval in the latent space may yield superior results.
-We will explore this in the next step.
+We will explore this in future work.
+A.6 Alternative Sampling Strategy: Importance Sampling
+In addition to the similarity-based sampling strategy with a fixed range $K$ described in Section˜3.3, we also explored the importance sampling approach.
+Instead of selecting auxiliary samples from a fixed-size neighborhood, this method assigns sampling probabilities to all training instances based on their similarity scores with the input function.
+Specifically, given an input function $a_{i}$, the probability of selecting any training sample $a_{j}$ as the auxiliary sample is defined as:
+$P(a_{j}|a_{i})=\frac{\text{exp}(\text{sim}(a_{i},a_{j})/\tau)}{\sum_{k=1}^{N}\text{exp}(\text{sim}(a_{i},a_{k})/\tau)},$ (17)
+where $\text{sim}(\cdot,\cdot)$ is the cosine similarity function defined in Equation˜6, $\tau$ is a temperature parameter controlling the sharpness of the distribution, and $N$ is the total number of training samples.
+We evaluated this importance sampling strategy on two representative problems: Darcy Flow with FNO and Heat Transfer with NORM. The results are presented in Table˜11.
+Table 11: Performance comparison of different sampling strategies
+Method Heat Transfer Darcy Flow
+Baseline (Direct Learning) 2.70e-4 3.70e-2
+DeltaPhi ($K=20$) 1.35e-4 3.31e-2
+DeltaPhi (Importance Sampling) 2.02e-4 3.22e-2
+The experimental results demonstrate that importance sampling can effectively enhance model performance, achieving 25.04% and 12.95% relative improvement on Heat Transfer and Darcy Flow respectively.
+Notably, on Darcy Flow, importance sampling even outperforms the fixed-range sampling strategy ($K=20$).
+This suggests that adaptive probability-based sampling could be a promising alternative to the fixed-range approach, particularly for problems where the similarity structure of the solution space is more complex.
+A.7 Comparison of Interpolation Methods for Cross-resolution Residual Learning
+In our framework, we employ Fourier-based interpolation for cross-resolution alignment and integration (as discussed in Section˜3.3 and Section˜3.4).
+This choice is advantageous as zero-padding in the frequency domain, which is equivalent to ideal sinc interpolation in the spatial domain, effectively upsamples the signal while exactly preserving the original low-frequency components without distortion.
+This characteristic aligns with the core principles of FNO-like architectures for zero-shot resolution generalization, which rely on accurately capturing the low-frequency dynamics of the physical solution.
+Table 12: Comparison of different interpolation methods on Darcy Flow.
+Interpolation Method Relative Error
+Fourier Interpolation 6.91e-2
+Bilinear Interpolation 7.36e-2
+Nearest Interpolation 7.39e-2
+To validate this choice, we conducted a comparison with other common interpolation methods, specifically Bilinear and Nearest Interpolation.
+The experiment was performed on the Darcy Flow problem, training on a $85\times 85$ resolution and testing on a $421\times 421$ resolution.
+The results, presented in Table˜12, show that Fourier-based interpolation achieves the lowest error, confirming its suitability for zero-shot generalization tasks.
+A.8 Symmetry Analysis of Residual Operator
+A key property of our residual learning formulation is the inherent symmetry between paired samples.
+To investigate whether the network naturally captures this property without explicit enforcement, we analyze the symmetry behavior of trained residual neural operators.
+Symmetry Metric. We introduce a Symmetry Loss metric to quantify how well the learned operator respects the expected symmetry for paired samples $(a_{i},u_{i})$ and $(a_{k_{i}},u_{k_{i}})$:
+$\text{Symmetry Loss}=\frac{\|\mathcal{G}^{\Delta}_{\theta}(a_{i},a_{k_{i}})+\mathcal{G}^{\Delta}_{\theta}(a_{k_{i}},a_{i})\|_{2}}{\|\mathcal{G}^{\Delta}_{\theta}(a_{i},a_{k_{i}})\|_{2}}$ (18)
+This metric measures the normalized discrepancy between residual predictions when the input order is reversed.
+A perfectly symmetric operator would yield a value of zero, while larger values indicate asymmetric behavior.
+Experimental Setup.
+We track this metric throughout the training process for our DeltaPhi-FNO model on the Darcy Flow problem.
+Importantly, we do not explicitly optimize for symmetry in the training objective; we simply observe how the standard training process affects symmetry properties.
+Table 13: Evolution of Symmetry Loss during training of DeltaPhi-FNO on Darcy Flow
+Metric Epoch 0 Epoch 300 Epoch 600 Epoch 900 Epoch 1200 Epoch 1500 Epoch 1800
+Symmetry Loss 2.01e+0 1.16e-1 8.00e-2 4.52e-2 4.12e-2 3.84e-2 3.22e-2
+Results and Analysis.
+The results in Table˜13 demonstrate that the model progressively learns to respect the inherent symmetry of the residual mapping without explicit guidance.
+Starting from a highly asymmetric state (Symmetry Loss of 2.01), the model naturally converges toward increasingly symmetric behavior, reaching a much lower Symmetry Loss of 3.22e-2 by the end of training.
+This emergence of symmetry-aware behavior validates that our residual learning approach inherently captures physically meaningful properties of the underlying systems.
+The network effectively learns that the relationship between two physical states should maintain consistent properties regardless of which state is considered the reference, aligning with fundamental principles of physical systems.
+These findings suggest potential future improvements through explicit symmetry enforcement during training or through specialized architectures designed to leverage this property, which could further enhance the physical consistency and sample efficiency of neural operator learning.
+A.9 Integration with Other Base Models
+The formulated residual mapping is a general operator learning task, fundamentally independent of the specific base architecture.
+Consequently, any neural operator can be employed to learn this mapping.
+In Table˜14, we include two more base models, Transolver [38] and HPM [25], on the Irregular Darcy problem.
+Table 14: Performance of DeltaPhi integrated with more base architectures.
+Model Relative Error
+Transolver (Direct) [38] 8.54e-3
+Transolver (DeltaPhi) 8.18e-3
+HPM (Direct) [25] 7.39e-3
+HPM (DeltaPhi) 6.67e-3
+The results show that DeltaPhi consistently improves the base models. In the future, it is significant to integrate DeltaPhi with more advanced architectures.
 Appendix B Experimental Detail
@@
 The low resolution data e.g.
-$22\times 2$ are obtained via uniformly downsampling operation.
+$22\times 22$ are obtained via uniformly downsampling operation.
 Navier-Stokes.
@@
 We use the public dataset from  [2].
-The viscosity of trajectories is $51e-$ 5 ($Re=2000$).
+The viscosity of trajectories is 1e51e-5 ($Re=2000$).
 The spatial resolution is $64\times 64$.
@@
 B.1.2 Irregular Domain
-For these irregular domain problems, we take the exactly same setting with NORM, more details about the dataset can be found in [10].
+For these irregular domain problems, we take the exact same setting with NORM, more details about the dataset can be found in [9].
 Irregular Darcy.
@@
 We employ the official implementation for the model architecture.
-The signature is set as $(-1,-1)$ and we pad lacked physical channels with zero.
+The signature is set as (1,1)(-1,-1) and we pad lacked physical channels with zero.
 All hyperparameters except the channel width (32 in CFNO) remain consistent with FNO.
@@
 B.3 Implementation Details
-Except for specific statements, the experimental details are as follows:
-  • Retrieval: For training, the auxiliary trajectory retrieval range $K$ is set as 20. During inference, the auxiliary trajectory with the most similarity with $a_{i}$ is retrieved.
-  • Input: The customized input function is partial (last 3 channels) $a_{k_{i}}$ and complete $u_{k_{i}}$.
-  • Optimizing: We use the Adam [46] optimizer with initial learning rate $0.001$ and weight decay $0.0001$.
+Except for specific statements, the experimental details are as follows: (a) Retrieval: For training, the auxiliary sample retrieval range $K$ is set as 20.
+During inference, the auxiliary sample with the most similarity with $a_{i}$ is retrieved.
+(b) Input: The customized input function is partial (last 3 channels) $a_{k_{i}}$ and complete $u_{k_{i}}$.
+(c) Optimizing: We use the Adam [57] optimizer with initial learning rate $0.001$ and weight decay $0.0001$.
 The StepLR scheduler with step size 100 and $\gamma=0.5$ is used to control the learning rate.
@@
 $L2={1\over N}\sum_{i=1}^{N}{\|\hat{u}_{i}-u_{i}\|_{2}\over\|u_{i}\|_{2}},$ (19)
-where $\hat{u}_{i}$ and $u_{i}$ are predicted trajectory and real trajectory respectively.
+where $\hat{u}_{i}$ and $u_{i}$ are predicted solution and real solution respectively.
 Appendix C Additional Method Detail
@@
 Given the training set (noted as $\{a^{train}_{i}\}_{i=1}^{N_{train}}$), suppose the input functions (noted as ${\{a^{test}_{i}\}}_{i=1}^{N_{test}}$) of pending solved physical trajectories is available.
-The initial value of $K$ could be calculated with the following steps: (1) Retrieve the most similar trajectory (noted as $\{a^{test}_{k_{i}}\}_{i=1}^{N_{test}}$) from the training set for every pending solving function.
+The initial value of $K$ could be calculated with the following steps: (1) Retrieve the most similar sample (noted as $\{a^{test}_{k_{i}}\}_{i=1}^{N_{test}}$) from the training set for every pending solving function.
 We denote the similarity values between $a^{test}_{i}$ and $a^{test}_{k_{i}}$ as $s^{test}_{i}$.
 (2) Calculate the maximum value $s^{test}_{max}$ of $\{s^{test}_{i}\}_{i=1}^{N}$.
-(3) For every training trajectory $a^{train}_{i}$, calculate its similarity with other training trajectories.
-Denote the index of $r$-th similar trajectory with $a^{train}_{i}$ as $k^{r}_{i}$, and the similar score between $a^{train}_{i}$ and $a^{train}_{k^{r}_{i}}$ as $s^{r}_{i}$.
+(3) For every training sample $a^{train}_{i}$, calculate its similarity with other training trajectories.
+Denote the index of $r$-th similar sample with $a^{train}_{i}$ as $k^{r}_{i}$, and the similar score between $a^{train}_{i}$ and $a^{train}_{k^{r}_{i}}$ as $s^{r}_{i}$.
 (4) The $K$ is calculated as the minimum $r$ satisfying $s^{test}_{max}\leq s_{i}^{r}$ for any $i\in[1,N^{train}]$.
-C.2 Variant Resolution Auxiliary Trajectory Integration
-Sometimes the auxiliary function $a_{k_{i}}$ has a different resolution.
-For example, when we need to predict high-resolution physical trajectory using low-resolution trained models, the auxiliary samples $a_{k_{i}}$ in the training set are more coarse-grained than the queried function $a^{H}$.
-In this scenario, we could interpolate the low-resolution auxiliary sample to a fine-grained grid, then concatenate the interpolated high-resolution grid with the queried function $a^{H}_{i}$.
-$a_{input}=\mathcal{F}^{inter}(a_{k_{i}})\oplus a^{H}.$ (13)
-In this work, we implement the interpolation method $\mathcal{F}^{inter}$ based on the Fourier Transform.
-Specifically, we first transform the coarse-grained input $u_{k_{i}}$ into Fourier Space, fill zeros in the high-frequency region to target resolution, and finally transform it back to spatial space.
-Our experimental results show the trained residual neural operator could generalize well in this interpolation scenario.
-C.3 Variant Resolution Trajectory Retrieval
-For the high resolution inference scenario, we could first downsample the high-resolution queried function $a^{H}$ to low resolution $a^{H}_{low}$, then calculate the similarity score:
-$\mathcal{F}^{ret}(a^{H},\mathcal{T})$ $=\text{argmax}_{j}\,\text{Sim}(a^{H}_{low},a_{j}),$ (14)
-$a^{H}_{low}$ $=\mathcal{F}^{low}(a^{H}),$
-where $\mathcal{F}^{low}$ is down sampling function. In this work, we implement $\mathcal{F}^{low}$ by erasing additional information in high-frequency regions based on the Fourier Transform.
-C.4 Complexity Analysis of Physical Trajectory Retrieval
-The computational cost for retrieval is negligible compared to neural network inference. Here is a detailed analysis of the computational cost for physical trajectory retrieval.
-The computational expensive is mainly related to two factors: the training set size (noted as $N_{train}$) and the number of physical field points (ie.
-the resolution of physical field) (noted as $N_{field}$).
-The computing process consists of the following two steps:
+C.2 Complexity Analysis of Auxiliary Sample Retrieval
+The computational cost for retrieval is negligible compared to neural network inference. Here we provide both theoretical complexity analysis and comprehensive experimental validation.
+Theoretical Analysis.
+The computational complexity primarily depends on two factors: the training set size (denoted as $N_{train}$) and the number of physical field points (i.e., the resolution of physical field) (denoted as $N_{field}$).
+The retrieval process consists of two main steps:
   • Step 1: Similarity Score Calculation.
-For every sample in the retrieval set, the similarity metric function (eg.
-Cosine Similarity) could be calculated with linear time complexity $O(N_{field})$.
-Thus the time complexity of calculating similarity scores for all retrieved samples is $O(N_{field}\cdot N_{train})$.
-  • Step 2: Ranking based on similarity scores.
-We use the simple sorting algorithm (Quicksort) to rank all samples based on their similarity score.
-The time complexity is $O(N_{train}\cdot\log N_{train})$.
-The overall time complexity for retrieval is $O(N_{field}\cdot N_{train}+N_{train}\cdot\log N_{train})$.
-Commonly, the numerical ranges for $N_{field}$ and $N_{train}$ are approximately between $10^{3}$ to $10^{4}$, and $10^{2}$ to $10^{4}$, respectively.
-Therefore, the real computational amount is about $O(10^{8})$, this could be calculated on most computers.
-In our implementation, the average retrieval time on Darcy Flow ($85\times 85$ resolution) is 2.67e-3 to 1.84e-2 seconds for different retrieval set sizes (100 to 900).
+For each sample in the retrieval set, computing the similarity metric (e.g., Cosine Similarity) has linear time complexity $O(N_{field})$.
+Thus, calculating similarity scores for all retrieved samples has complexity $O(N_{field}\cdot N_{train})$.
+  • Step 2: Similarity Score Ranking. Using QuickSort to rank samples based on similarity scores has complexity $O(N_{train}\cdot\log N_{train})$.
+The overall time complexity is $O(N_{field}\cdot N_{train}+N_{train}\cdot\log N_{train})$.
+With typical values of $N_{field}$ and $N_{train}$ ranging from $10^{3}$ to $10^{4}$ and $10^{2}$ to $10^{4}$ respectively, the computation is feasible on standard hardware.
+Experimental Validation. We conducted extensive experiments to measure the actual computational overhead of retrieval:
+  • Retrieval Time Analysis: We tested retrieval time across different training set sizes on both CPU and GPU (single RTX 3090) for Darcy Flow ($85\times 85$ resolution):
+    Table 15: Retrieval time (in seconds) for different training set sizes
+    Device 100 300 500 700 900
+    CPU 2.67e-3 5.57e-3 9.05e-3 1.32e-2 1.66e-2
+    GPU 1.89e-4 1.83e-4 1.64e-4 2.43e-4 3.06e-4
+  • End-to-End Inference Time: We compared the total inference time per example between direct learning and residual learning across different models:
+    Table 16: Average inference time (in seconds) comparison
+    Method FNO FFNO Galerkin
+    Direct Learning 1.20e-3 2.39e-3 5.01e-3
+    Residual Learning 1.67e-3 2.85e-3 5.58e-3
+    Increased Time 4.67e-4 4.66e-4 5.75e-4
+The experimental results demonstrate that:
+  • GPU-based retrieval is highly efficient, taking only 0.2-0.3ms across different training set sizes.
+  • The total additional time for residual learning, including retrieval and data preparation, is consistently around 0.5ms.
+These comprehensive experiments confirm that the auxiliary sample retrieval introduces negligible computational overhead and does not significantly impact the practical deployment of our method.
+C.3 Discussion on Boundary Conditions
+In this work, DeltaPhi does not explicitly process boundary conditions as a separate input.
+Instead, it implicitly handles them by leveraging the full auxiliary solution $u_{k_{i}}$ as a strong physical prior.
+This auxiliary solution, which is a key input to the model, already satisfies the boundary conditions of its corresponding system.
+The model’s effectiveness in handling complex boundary effects is demonstrated by its strong performance on irregular domain problems, such as Pipe Turbulence and Blood Flow, which feature complex inlet/outlet conditions.
+While our current approach is effective, future work could explore enhancing the retrieval mechanism with a boundary-aware similarity metric for enhanced residual neural operator.
 Appendix D Visualization Analysis Detail
 D.1 Similarity between $u(x)$ Visulization Detail
-We visualize the correlation between $u(x)$ normalized distance and $a(x)$ similarity rank (shown in Figure 4) as following steps: Firstly, taking the training set $\mathcal{T}$ as retrieval set, we sample auxiliary trajectory $(a_{k_{i,r}},u_{k_{i,r}})$ with $r$-th top similarity score for every test sample $(a^{test}_{i},u^{test}_{i})$ in the test set.
+We visualize the correlation between $u(x)$ normalized distance and $a(x)$ similarity rank (shown in Figure˜4) as following steps: Firstly, taking the training set $\mathcal{T}$ as retrieval set, we retrieve auxiliary sample $(a_{k_{i,r}},u_{k_{i,r}})$ with $r$-th top similarity score for every test sample $(a^{test}_{i},u^{test}_{i})$ in the test set.
 Here $r$ represents the similarity rank (higher $r$ indicates lower similarity), and we take gradually increasing values from $1$ to $80$ for $r$.
 Next, we calculate the normalized distance between $u^{test}_{i}$ with its auxiliary sample $u_{k_{i,r}}$.
-The normalized distance is defined the same as the relative error Equation 14.
+The normalized distance is defined the same as the relative error in Equation˜19.
 D.2 Label Distribution Visualization Detail
-We visualize the label distribution (as shown in  Figure 5) by reducing the high-dimension function fields to 2 dimensions utilizing Principle Component Analysis.
-Specifically, for both direct operator mapping and residual operator mapping, we first optimize a PCA model on the training labels (100 labels) and then calculate the dimension-reduced labels for every trajectory in the training and testing set.
+We visualize the label distribution (as shown in  Figure˜5) by reducing the high-dimension function fields to 2 dimensions utilizing Principle Component Analysis.
+Specifically, for both direct operator mapping and residual operator mapping, we first optimize a PCA model on the training labels (100 labels) and then calculate the dimension-reduced labels for every sample in the training and testing set.
 Finally, we visualize every two-dimension label point on the graph as well as the ellipses representing the stand deviation and the range along each dimension.